### v2

# Demystifying Language Model Forgetting with Low-rank Example Associations

Abstract
Large Language models (LLMs) suffer from forgetting of upstream data when fine-tuned. Despite efforts on mitigating forgetting, few have investigated whether, and how forgotten upstream examples are dependent on and associated with newly learned tasks. Insights on such associations enable efficient and targeted mitigation of forgetting. In this paper, we empirically analyze forgetting (measured in log-perplexity increase) that occurs in $N$ upstream examples of language modeling or instruction-tuning after fine-tuning LLMs on one of $M$ new tasks, visualized in $M\times N$ matrices. We demonstrate that the matrices display simple low-rank patterns, often well-approximated with multiplicative scalar effects of upstream examples and newly learned tasks. We also examine fine-grained associations with visualization and statistics. Leveraging the low-rank nature of the associations, we predict forgetting of upstream examples when fine-tuning on unseen tasks with matrix completion over the empirical associations. This enables fast identification of most forgotten examples without expensive inference on the entire upstream data. The approach, despite simplicity, outperforms prior approaches that learn semantic relationships of learned tasks and upstream examples with LMs for predicting forgetting. We demonstrate the practical utility of our analysis by showing statistically significantly reduced forgetting as we upweight predicted examples for replay at fine-tuning. Project page: https://inklab.usc.edu/lm-forgetting-prediction/
1 Introduction
There has been a growing need for long-term usability of LLMs. Although fine-tuning allows efficient and incremental adaptation of models, it risks catastrophic forgetting (McCloskey & Cohen, 1989; Goodfellow et al., 2014) of upstream knowledge learned in the pre-training or instruction-tuning phase, causing unintended prediction changes over known information. This is problematic for the stability of online deployed LLM systems, limiting the feasibility of continual fine-tuning in practice (Raffel, 2023; Shi et al., 2024).
While extensive works have developed algorithms to mitigate forgetting (Wu et al., 2024), the increasing scale of the model and data necessitates more efficient and targeted approaches. A key aspect of the challenge is understanding when and what LLMs forget, prompting analysis into the patterns of frequently forgotten examples (Toneva et al., 2019; Maini et al., 2022; Zhang & Wu, 2024), and the impacts of models and hyperparameters on forgetting (Mirzadeh et al., 2022; Kalajdzievski, 2024; Que et al., 2024; Ibrahim et al., 2024). However, how the associations between learned tasks and upstream examples inform forgetting remains under-explored. Understanding such associations creates potential for targeted mitigation of forgetting over specific upstream examples when a new task is learned. While theoretical and empirical study indicate associations between learned and forgotten tasks in shallower neural networks (Lee et al., 2021; Goldfarb et al., 2024; Ramasesh et al., 2021), the associations are under-explored for LLMs, or measured regarding upstream data of language modeling or instruction-tuning.
In this paper, we empirically study the associations between learned tasks and forgotten upstream examples of language modeling or instruction-tuning. We experiment with OLMo-1B, OLMo-7B, OLMo-7B-Instruct (Groeneveld et al., 2024) and MPT-7B (Team, 2023) models where upstream data is open-source. We perform statistics of forgetting (in log perplexity increase) over $N$ upstream examples, after fine-tuning the model on one of $M$ unseen instruction-tuning tasks, and represent the statistics in a $M\times N$ matrix. Afterwards, we visualize the matrices and fit the observations with statistical models to analyze the associations. Figure 1 illustrates the analysis setup.
Our findings suggest that the associations between learned tasks and forgotten examples are usually simple and display a low-rank pattern, as simple regression models can decently fit of the associations. We extract more fine-grained associations with visualization and statistics, showing cases where specific upstream examples are forgotten when learning a new task. We demonstrate that forgetting does not correlate with many predefined similarity metrics of learned tasks and upstream examples, such as token overlap or gradient dot-products, making these metrics poor predictors of forgetting. Inspired by the findings, we directly utilize statistics of forgetting to predict example forgetting on unseen tasks by solving a matrix completion problem over the association matrices, analogical to collaborative filtering (Sarwar et al., 2001) in recommender systems, achieving both efficiency and interpretability. Our $k$-nearest neighbor (KNN) model outperforms prior approaches that learn semantic relations of two examples with LMs (Jin & Ren, 2024) . We verify the benefit of prediction by upweighting examples with higher predicted forgetting during replay as we fine-tune LLMs on new instruction-tuning tasks, achieving statistically significant improvement in alleviating forgetting compared to replaying random examples.
Figure 1: The problem setup of analyzing the associations between learned tasks and forgotten upstream examples as we fine-tune LLMs on one of unseen new tasks. Over total $N$ upstream examples and $M$ unseen tasks, we measure and record forgetting in a $M\times N$ matrix and analyze the statistical properties of the associations.
To summarize, the contributions of this paper are (1) an empirical analysis on how forgotten examples are associated with learned tasks in representative 1B and 7B language models, and (2) a novel approach of predicting example forgetting by solving a matrix completion problem over the empirical associations, and (3) a practical and efficient algorithm to mitigate forgetting during LLM fine-tuning by upweighting upstream examples for replay according to the predicted forgetting.
2 Problem and Analysis Setup
In this section, we start by formally defining the metrics of forgetting and set up the problem formulation of analyzing the associations between learned tasks and forgotten examples. We then introduce models and datasets used for collecting the statistics.
2.1 Collecting Statistics of Forgetting
Upstream examples and learned tasks. Large Language models (LLMs) are commonly pre-trained with language modeling objectives over a massive collection for corpora, with some LLMs further post-trained (instruction-tuned) to better follow human instructions. We use upstream data to refer to language modeling or instruction tuning training data used at the pre-training or post-training phase of LLMs. For upstream data of language modeling, we define each upstream example $x_{j}\in x_{1..N}$ as a chunk of document (e.g., a Wikipedia article) of a model-specific maximum number of tokens. For instruction tuning, each $x_{j}\in x_{1..N}$ corresponds to a pair of instructions and ground truth responses.
Measuring forgetting. We fine-tune an LLM (or an instruction-tuned LLM) on one unseen instruction-tuning task $T_{i}$ from a collection of tasks $T_{1..M}$. This results in $M$ separately fine-tuned models $f_{1..M}$. We then evaluate performance degradation on each upstream example $x_{i}\in x_{1..N}$. We measure log perplexity as the performance metric as they are applicable to both language modeling and instruction tuning, and are known to correlate well with other dataset-specific metrics (Hoffmann et al., 2022). For language modeling, log perplexity of an example $x_{i}$ is averaged over all tokens in the example; for instruction tuning, we average log perplexity over the ground truth response tokens only. We measure forgetting $z_{ij}$ that occurs on an upstream example $x_{j}\in x_{1..N}$ as increase (degradation) in log perplexity after fine-tuning the LM on a new task $T_{i}\in T_{1..M}$. We record forgetting $z_{ij}$ in an association matrix $Z\in\mathbb{R}^{M\times N}$.
2.2 Models and Datasets
Our analysis requires access to upstream data of LLMs for measuring their log perplexity changes occurred after fine-tuning. We experiment with OLMo-1B, OLMo-7B, OLMo-7B-Instruct, and MPT-7B where upstream data of language modeling and instruction-tuning is open-source.
OLMo-1B and 7B. OLMo models are pre-trained on Dolma (Soldaini et al., 2024), a massive collection of text covering diverse domains such as Wikipedia and news articles. We fine-tune LMs over 66 tasks from FLAN-V2 (Longpre et al., 2023), 11 tasks from Tulu V2 (Ivison et al., 2023), and 8 tasks from Dolly (Conover et al., 2023), obtaining 85 fine-tuned models. The collections cover diverse instruction-tuning tasks such as reading comprehension, math reasoning, and safety alignment. We then evaluate log perplexity increase on a 1% subset of Dolma-v1.6-Sample. Each upstream example is a maximum 2,048-token document from Dolma, resulting in 141,816 examples. For 7B models, We perform either full-parameter fine-tuning or LoRA adapter tuning (Hu et al., 2022) only, noted as OLMo-7B (Full FT) or OLMo-7B (LoRA) respectively.
OLMo-7B-Instruct. OLMo-7B-Instruct models are instruction-tuned on Tulu V2. In our experiments, we few-shot fine-tune OLMo-7B-Instruct over new task data from MMLU (Hendrycks et al., 2021), BBH (Suzgun et al., 2022), TruthfulQA (Lin et al., 2022), and Dolly, and evaluate log perplexity increase over a stratified sample of 10,718 examples from Tulu v2 as upstream examples. The number of training examples in these datasets are smaller (on the scale of tens to hundreds of examples), making the setup closer to LLM error editing (Yao et al., 2023; Zhang et al., 2024). We perform either full-parameter fine-tuning or LoRA fine-tuning.
MPT-7B. MPT models are pre-trained on a diverse collection of corpora. We fine-tune the LM over the same 85 instruction-tuning tasks as OLMo models and evaluate forgetting on Redpajama (Computer, 2023), involved as a part of the pretraining corpora of MPT. We sample 10,000 maximum 2,048 token documents from Redpajama as upstream examples.
We use a learning rate of $2e^{-6}$ for full-parameter fine-tuning and $1e^{-4}$ for LoRA fine-tuning. We include the other training details in Appendix A.
(a) OLMo-1B; forgetting over Dolma after full-parameter fine-tuning on FLAN and Tulu.
(b) OLMo-7B; forgetting over Dolma after full-parameter fine-tuning on FLAN and Tulu.
(c) OLMo-7B-Instruct; forgetting over Tulu after full-parameter fine-tuning on unseen instruction-tuning tasks.
Figure 2: Visualized matrices of associations between learned tasks and forgotten examples. We plot forgetting (log-perplexity increase) that occurs on an upstream example (in $x$-axis) after learning a new task (in $y$-axis). Log-perplexity increase can be zero or negative, which implies no forgetting.
3 Associations between Learned Tasks and Forgotten Examples
In this section, we address the research questions regarding the associations between learned tasks and forgotten upstream examples represented in the $M\times N$ association matrices $Z$. We start by examining how complicated are the associations represented by the matrices $Z$ (Sec. 3.1). We then extract and analyze more fine-grained associations in $Z$ (Sec. 3.2), and examine how the associations correlate with similarity measures of learned tasks and upstream examples (Sec.3.3).
3.1 How Complicated are the Associations?
In a hypothetical extreme case, the associations between the learned tasks and forgotten examples can be very simple, e.g., where upstream examples are forgotten regardless of learned tasks; or it can be highly complicated, e.g., where different new tasks cause very specific subsets of upstream examples to be forgotten. Figure 1 (right) illustrates how such simple or complicated associations may display in visualized matrices. To understand the actual associations, we start by visualizing the association matrix $Z$ collected in the setups described in Sec. 2.2, and perform quantitative analysis as we examine whether simple regression models can fit $Z$ with low error.
Analysis on the visualized association matrices. We visualize the association matrices $Z$ in Figure 2 for OLMo-1B, OLMo-7B, OLMo-7B-Instruct as we perform full-parameter fine-tuning. The matrices of MPT and LoRA fine-tuning of OLMo-7B and OLMo-7B-Instruct are presented in Figure 7 in Appendix. We see $Z$ generally displays a neat and simple pattern. We notice that certain upstream are more prone to forgetting, while some are never forgotten (displayed as all-zero columns). Nevertheless, in some cases, upstream examples that are never forgotten elsewhere are forgotten by learning specific new tasks (e.g. wizardlm examples are almost only forgotten after learning tasks from TruthfulQA in Figure 2(c)). It implies the association is a mixture of simple and more complicated patterns.
Quantitative evaluation of simplicity of the associations. We quantitatively measure how well the association matrices $Z$ can be approximated with simple regression models with a small number of learnable parameters. We consider (1) additive linear models, where $z_{ij}=b+\alpha_{i}+\beta_{j}+\epsilon$, where $\alpha_{i}$ and $\beta_{j}$ are learnable parameters associated with each new task or upstream example. (2) multiplicative models (SVD with rank $r$=1), where $z_{ij}=s\alpha_{i}\beta_{j}+\epsilon$. Both models involve $M+N$ learnable parameters plus a bias term. We then measure $R^{2}$ as the metrics of determining how well the regression models fit the association matrices $Z$. Let $f_{ij}$ be the fitted value, $R^{2}$ is defined as $1-\sum_{i,j}(z_{ij}-f_{ij})^{2}/\sum_{i,j}(z_{ij}-\bar{Z})^{2}$. $R^{2}$ ranges between 0 and 1, indicating the portion of total variance in $Z$ that is explained by the regression models. We report $R^{2}$ of additive and multiplicative models under different setups in Figure 3 (a).
Multiplicative models fit the associations better. In 5 out of 6 setups presented in Figure 3 (a), the multiplicative model achieves a better fit than additive models at the same number of trainable parameters. An interpretation of the multiplicative models is that each upstream example is associated with their tendency of being forgotten $(\beta_{j})$; the learned tasks trivially determine how fast all upstream examples are forgotten (with $\alpha_{i}$). This also creates cases where certain upstream are never forgotten $(\beta_{j}\approx 0)$, or learn certain tasks causing little forgetting on any upstream examples $(\alpha_{i}\approx 0)$.
(a) Fitting $Z$ with additive or multiplicative models
(b) Fitting $Z$ with progressively higher rank matrices
Figure 3: $R^{2}$ of fitting the associations between learned tasks and forgotten examples with (a) simple additive linear or multiplicative models, or (b) progressively more expressive models.
A large portion of the variance in the association can be explained by the multiplicative model. We notice the multiplicative models achieve $R^{2}$ between 0.43 to 0.72 in different setups despite their simplicity, suggesting generally simple associations between learned tasks and forgotten examples. This finding is emphasized by the diversity of the learned tasks we considered (spanning from reading comprehension, coding, to safety alignment) and the broad coverage of domains in upstream language modeling and instruction-tuning data. Nevertheless, we also notice the $R^{2}$ scores are relatively lower (0.4 to 0.5) on OLMo-7B and OLMo-7B-Instruct, implying more complicated associations than the other setups (OLMo-1B, MPT-7B, and LoRA fine-tuning of OLMo-7B models), which elicits our next research question.
3.2 What are More Fine-Grained Associations?
Figure 4: Patterns in associations captured by $k$-th component in SVD of the assoication matrix $Z$, exemplified with OLMo-7B-Instruct (full-parameter tuning, also in Fig. 2(c)). The highlighted regions display patterns that TruthfulQA tasks cause more forgetting ($k=2$) and MMLU causes more forgetting on certain tasks in FLAN v2 and Open-Orca ($k=3$).
We extract more fine-grained associations where fine-tuning on a task causes a specific set of upstream examples to be forgotten. We perform Singular Value Decomposition (SVD) over the association matrices $Z$ and progressively reconstruct $Z$ with up to $r$-th singular values and vectors as $Z_{r}=\sum_{k=1}^{r}s_{k}\bm{\alpha}_{k}\bm{\beta}_{k}^{T}$. $Z_{r}$ is also the optimal rank-$r$ matrix that minimizes the Frobenius norm $||Z-Z_{r}||_{F}$ and the $R^{2}$ score according to the property of SVD. The first component $Z_{1}$ corresponds to the multiplicative model we examined in Sec. 3.1; the rest of components $Z_{2..M}$ captures fine-grained associations in $Z$.
SVD reveals fine-grained associations. We present $R^{2}$ of fitting the assoication with $Z_{r}$ with progressively larger rank $r$ in Figure 3 (b). We notice that for OLMo-7B and OLMo-7B-Instruct (full-parameter fine-tuning), $R^{2}$ quickly increases to 0.69 and 0.78 with $Z_{3}$ (reconstruction with up to the third component). We visualize the matrices reconstructed from each component $s_{k}\bm{\alpha}_{k}\bm{\beta}_{k}^{T}$ in SVD of $Z$. Figure 4 provides an example of patterns captured by the $k$-th component. We demonstrate how the visualization extracts patterns of forgetting that is conditional on the learned tasks.
3.3 How do the Similarity Measures Interpret the Associations?
We examine whether the associations between learned tasks and forgotten examples are interpretable from the similarity between the learned tasks and upstream examples. We consider (1) heuristic similarity measures, such as textual similarity, and (2) theoretically grounded approximations, such as inner products of gradients.
Textual similarity. We measure textual similarity $z^{\textrm{text}}_{ij}$ between learned tasks and forgotten examples with TF-IDF vectorized over each pair of learned tasks $T_{i}$ and upstream examples $x_{j}$.
Inner products between projected gradients and model weight updates. The increase of the log perplexity ${z_{ij}}$ (also known as the cross-entropy loss) can be approximated with inner products $z^{\textrm{g-w}}_{ij}=\langle\nabla_{\theta}f(x_{j}),\theta_{T_{i}}-\theta_{0}\rangle$ under first-order Taylor expansion (Lee et al., 2019; Doan et al., 2020), where $\nabla_{\theta}f(x_{j})$ is the gradient of the loss of $x_{j}$ at the initial model before fine-tuning, and $\theta_{T_{i}}-\theta_{0}$ are the updates in the model weights after fine-tuning. Following Park et al. (2023); Xia et al. (2024), we use a random projection matrix $\bm{P}\sim\mathcal{N}_{|\theta|\times d}(0,1)$ to reduce the dimension of the gradients or the weight changes to save the cost of storing pre-computed statistics, which preserves the inner products with high probability (Johnson & Lindenstrauss, 1984).
Inner products between projected gradients. We also measure the negative inner products of the loss gradients between the upstream example $x_{j}$ and a learned task $T_{i}$, $z^{\textrm{g-g}}_{ij}=-\langle\nabla_{\theta}f(x_{j}),\nabla_{\theta}f(T_{i})\rangle$, as an approximation of forgetting (Lopez-Paz & Ranzato, 2017; Chaudhry et al., 2019).
Table 1: Spearman’s correlation coefficients between various measures of similarity and the actual forgetting on upstream examples after fine-tuning on a task.
Similarity Metrics Spearman’s $\rho$
Textual $z^{\textrm{text}}$ 0.004
Gradient-Weight Difference $z^{\textrm{g-w}}$ 0.005
Gradient-Gradient $z^{\textrm{g-g}}$0.007
Forgetting correlates poorly with similarity measures of learned tasks and upstream examples. We evaluate Spearman’s correlation between the actual forgetting $z_{ij}$ and the various similarity measures $\{z^{\textrm{text}}_{ij},z^{\textrm{g-w}}_{ij},z^{\textrm{g-g}}_{ij}\}$ on OLMo-1B and summarize the results in Table 1. We notice that none of the similarity measures correlates with the actual forgetting, with a correlation $\rho<0.01$. An interpretation of the low correlation is that the model weights deviate from the initial weights after fine-tuning to a region where first-order approximation of forgetting does not hold. We further visualize the matrices of $z^{\textrm{g-w}}_{ij}$ and $z^{\textrm{g-g}}_{ij}$ in Fig. 8 in Appendix and provide a side-by-side comparison with the matrices of forgetting $Z$. The visualization displays distinct patterns among the three matrices. These results imply that although the association matrices $Z$ display a simple pattern, they are not well-interpreted with common similarity measures of learned tasks and forgotten examples.
4 Predicting Example Forgetting with Association Matrix Completion
We utilize our findings in Sec. 3 to predict example forgetting as the model learns a new task, a problem also studied in prior works (Jin & Ren, 2024). Finding out the most forgotten examples allows better spot of the behavior changes of the models, enabling efficient and targeted approaches to mitigate forgetting, e.g., by replaying these examples (Aljundi et al., 2019a; Wang et al., 2024). Although the ground truth forgetting can be directly obtained by running inference with the fine-tuned model over the upstream data, this requires extensive computational resources. We restrict the prediction methods to be computationally efficient.
Our analysis in Sec. 3 suggests that (1) the associations between learned tasks and forgotten examples display simple statistical patterns, while (2) the associations correlate poorly with many similarity metrics. Therefore, we hypothesize that leveraging the statistics of forgetting allows more effective prediction of forgetting than leveraging the contents of the tasks and examples. Following this intuition, we formulate prediction of example forgetting as a matrix completion problem over the empirical associations $Z$, analogical to collaborative filtering in recommender systems (Sarwar et al., 2001), where scalable approaches are studied extensively. We start by setting up the problem formulation of predicting example forgetting, and evaluate the performance of prediction of different approaches. We present reduced forgetting by utilizing the prediction outcomes during fine-tuning.
4.1 Training and Evaluation of Forgetting Prediction
Our goal is to accurately predict forgetting $z_{ij}$ over upstream examples $x_{1..M}$ when the model is fine-tuned on an unseen task $T_{j}$ with a prediction model $g$, without running expensive LLM inference on all $x_{1..M}$. To evaluate this, we create training and test splits by partitioning the set of fine-tuning tasks (noted as $\mathcal{T}_{\textrm{train}}$ and $\mathcal{T}_{\textrm{test}}$) and the rows of the association matrices $Z$. We further control whether the $\mathcal{T}_{\textrm{train}}$ and $\mathcal{T}_{\textrm{test}}$ are from the same category of the tasks to test both in-domain and out-of-domain generalization ability of the prediction models. For OLMo-1B and 7B experiments, we use FLAN and the in-domain tasks and Tulu and Dolly as out-of-domain testing tasks. For OLMo-7B-Instruct experiments, we use MMLU and BBH as in-domain tasks and use TruthfulQA and Dolly as out-of-domain testing tasks. Details about the tasks included in the training, in-domain testing, and out-of-domain testing sets are discussed in Appendix B.
Figure 5: The training and testing setup of predicting example forgetting with association matrix completion.
To apply matrix completion for predicting forgetting, a few entries $z_{ij}$ should be known when a new fine-tuning task $T_{i}\in\mathcal{T}_{\textrm{test}}$ (row $i$) is introduced. We therefore assume access to the ground truth forgetting $z_{ij}$ of a tiny random set $\mathcal{S}$ ($|\mathcal{S}|=30$) of upstream examples for $T_{i}\in\mathcal{T}_{\textrm{test}}$, noted as seed forgetting $z_{i}^{\mathcal{S}}=\{z_{ij}|x_{j}\in\mathcal{S}\}$. Obtaining seed forgetting typically takes only a few seconds by running inference with the model fine-tuned on $T_{i}$ over $\mathcal{S}$; we then predict forgetting of the rest $10k-100k$ upstream examples. Figure 5 illustrates an example of the train-test partition, seed forgetting, and the forgetting to be predicted. We use Root Mean Squared Error (RMSE) over the $\mathcal{T}_{\textrm{test}}$ as the metrics of predicting example forgetting.
Matrix completion approaches. We run matrix completion algorithms including additive linear, SVD, and k-nearest neighbors (KNN) models. The additive linear and the SVD models are introduced earlier in Sec. 3. Given the seed forgetting $z_{i}^{\mathcal{S}}$ of a task $T_{i}\in\mathcal{T}_{\textrm{test}}$, KNN finds tasks from $\mathcal{T}_{\textrm{train}}$ that have similar patterns of forgetting over the seed upstream examples $\mathcal{S}$. KNN computes an average of forgetting of top-k similar tasks from $\mathcal{T}_{\textrm{train}}$ weighted by their similarity as the prediction of forgetting caused by $T_{i}\in\mathcal{T}_{\textrm{test}}$ on the upstream examples $x_{1..M}$.
Comparators of predicting forgetting. We compare with a prior approach by Jin & Ren (2024) that leverages learned similarity between learned tasks and upstream examples by a trainable LM to predict forgetting. This prior work, however, focuses on predicting forgetting while fixing one single error in LM predictions; we extend the approach to predicting forgetting after fine-tuning models over a task (i.e., a set of examples). The extended approach encodes upstream examples $x_{j}$ and the training examples of the learned task $x_{i}^{1..N_{i}}\in T_{i}$ with a trainable LM encoder $h(\cdot)$ to obtain their representations. The final prediction is made with a regression head over the inner products of two representations $\langle h(x_{j}),\frac{1}{N_{i}}\sum_{N_{i}}h(x_{i})\rangle$.
We leave the implementation details of matrix completion approaches and the learned similarity approach in Appendix B.
Table 2: RMSE of predicting example forgetting over a held-out set of upstream examples after fine-tuning LMs on unseen new tasks. We report averaged performance over different seed sets ($\mathcal{S}$) of upstream examples with known ground truth forgetting beforehand.
In-Domain Out-of-Domain
OLMo-1B OLMo-7B MPT-7B OLMo-7B-Inst. OLMo-1B OLMo-7B MPT OLMo-7B-Inst.
Full FT Full FT LoRA Full FT Full FT LoRA Full FT Full FT LoRA Full FT Full FT LoRA
Additive 2.81 7.40 3.50 13.33 15.57 6.12 2.81 5.83 7.01 10.02 38.90 21.22
SVD 2.88 7.53 3.49 12.41 14.98 6.24 2.86 5.84 7.16 8.27 39.03 21.26
KNN 2.79 7.33 3.45 12.80 14.30 5.54 2.84 5.83 6.83 7.71 38.77 20.82
Similarity 3.84 9.29 5.45 14.00 16.23 6.19 3.93 7.64 8.52 10.97 42.38 23.47
4.2 Mitigating Forgetting with Predicted Forgetting
Leveraging predicted forgetting for mitigating forgetting. We examine the practical utility of predicting forgetting as we sparsely replay upstream examples during forgetting following Jin & Ren (2024). Sparse replay of upstream examples is known as an effective and model-agnostic way to mitigate forgetting (de Masson D’Autume et al., 2019; Ibrahim et al., 2024). We replay one mini-batch of upstream examples every 32 training steps while fine-tuning on a new task. We perform targeted mitigation of forgetting by prioritizing examples that are predicted to suffer more from forgetting. This is achieved with weighted sampling of upstream examples $x_{j}$ proportional to $\exp{(\hat{z}_{ij}/\tau)}$, where $\hat{z}_{ij}$ are the predicted forgetting and $\tau$ is a temperature hyperparameter set as 0.1.
Baselines of mitigating forgetting. We compare with diverse strategies of selecting upstream examples for sparse replay. As a main comparator, we examine whether weighted sampling with the predicted forgetting improves over random sampling of upstream examples (Random). We compare with Maximally Interfered Retrieval (MIR) (Aljundi et al., 2019a), a selection strategy sharing the similar notion of importance that forgotten examples should be selected for replay. The approach performs bi-level sampling by selecting the most forgotten examples from a small random subset of upstream data (set as $|\mathcal{S}|=30$). In addition, we apply strategies that consider different definitions of upstream example importance. We examine an approach based on perplexity thresholds (PPL) (Marion et al., 2023), which samples upstream data of which the perplexity is around the median of the distribution. For OLMo-1B, we also sample replayed examples proportional to the gradient inner products (Grad-Prod) we evaluated in Sec. 3.3, in a similar vein to coreset selection approaches that utilize gradient information (Park et al., 2023; Xia et al., 2024).
Metrics. We measure log-perplexity increase over a held-out subset of 10,000 examples from the upstream data. This ensures none of the test examples are selected for replay by any of the example selection strategies.
4.3 Results of Predicting and Mitigating Forgetting
Results of predicting example forgetting. Table 2 summarizes the error of predicting example forgetting over tasks from the in-domain and out-of-domain test splits. We see matrix completion approaches consistently outperform the learned similarity model in the prior work. Among the three matrix completion approaches, we notice that KNN models in general achieve the lowest prediction error. Given this finding, we primarily leverage predictions given by the KNN model to reduce forgetting during fine-tuning.
(a) 1B, FLAN
(b) 1B, Tulu&Dolly
(c) 7B, FLAN
(d) 7B, Tulu&Dolly
(e) 7B-Inst, Dolly
Figure 6: Log perplexity over upstream data as replay upstream examples selected by different approaches after fine-tuning OLMo or OLMo-Instruct models over in-domain or out-of-domain test tasks. The solid horizontal lines indicate the log perplexity before fine-tuning (i.e., no forgetting). The dash lines show the log perplexity achieved by upweighting upstream examples according to their ground truth forgetting. * and ** indicate statistical significance of improvement ($p<0.05$ or $p<0.005$) compared to replaying random examples in paired $t$-tests over all fine-tuning tasks.
Mitigating forgetting with the predicted forgetting. We leverage the forgetting predicted by the KNN model and upweight examples during replay following our procedure in Sec. 4.2. Figure 6 summarizes log perplexity after fine-tuning over the held-out (never replayed) upstream data as we apply different upstream example selection approaches. We visualize the relative comparisons of these approaches to no forgetting of upstream examples (log perplexity before fine-tuning). We notice that example selection based on gradient inner products (Grad) or perplexity threshold (PPL), mainly applied for identifying important training data for a task in prior works, does not show improvement in mitigating forgetting compared to replaying random examples. This implies that the notions of example importance in these works are different from how easily the examples are forgotten. We also notice that MIR does not improve over random sampling in our setup, likely because of the small size of the retrieval candidates relative to the upstream examples. We see upweighting examples with ground truth forgetting (GT) consistently reduces forgetting compared to random examples. Replaying with KNN-predicted forgetting also consistently reduces forgetting, achieving statistical significance in 4 out of 5 tested setups.
Limitations of replaying with forgetting prediction. We note that our current practice of replaying with predicted forgetting requires two runs of fine-tuning: a replay-free run of fine-tuning after which the seed forgetting will be evaluated, and another run while replaying examples with the predicted forgetting. This creates computational overhead equivalent to one extra run of fine-tuning, which is identical to the MIR baseline. Nevertheless, the practice is still efficient when the training set of fine-tuning is much smaller than the upstream data. Future works can predict forgetting on-the-fly during fine-tuning, mitigating the overhead of fine-tuning the model twice.
5 Related Works
Factors that affect forgetting. In this paper, we primarily studied how the associations between learned and forgotten examples inform forgetting. Prior works have studied various factors that affect forgetting of the models, such as (1) type and size of the LM (Mehta et al., 2021; Scialom et al., 2022; Kalajdzievski, 2024; Mirzadeh et al., 2022) (2) trainable parts of the model (e.g., LoRA, soft prompts, or full-model tuning) (Biderman et al., 2024a; Razdaibiedina et al., 2023) (3) hyperparameters such as learning rate (Ibrahim et al., 2024; Winata et al., 2023), dropout (Goodfellow et al., 2014), number of training steps (Biderman et al., 2024b; Kleiman et al., 2023) (4) optimizer (Lesort et al., 2023) and training algorithms (e.g., various continual learning algorithms) (Shi et al., 2024; Wu et al., 2024), (5) the upstream examples or the knowledge themselves (Toneva et al., 2019; Zhang & Wu, 2024). Future works can study how various factors affect the associations between learned tasks and forgotten examples. We consider empirical and theoretical study on the effect of task similarity on forgetting to be most relevant to ours.  Ostapenko et al. (2022) empirically study relationships between task similarity and forgetting in foundation models over a sequence of newly learned tasks; our work instead focuses on forgetting of upstream data of LLMs. Theoretical study by Doan et al. (2020); Ding et al. (2024); Evron et al. (2022) dissects effects of the learned tasks on forgetting in linear models or around model initialization. We believe research on interpretations of forgetting (Tao et al., 2023; Zhao et al., 2023; Kotha et al., 2024) is complementary to ours and can potentially explain in the future why the associations in $Z$ are often simple, and in which circumstances the associations become more complicated.
Data selection and data attribution. Related to our work, data attribution studies faithful algorithms to find training examples that account for a prediction (Koh & Liang, 2017; Ilyas et al., 2022) from a pool of training examples. Park et al. (2023); Xia et al. (2024); Li et al. (2024); Liu et al. (2024) study the problem of selecting a subset of training data that maximizes performance on a given domain or task at a fixed budget for LLMs. Feldman & Zhang (2020); Tirumala et al. (2022); Biderman et al. (2024b); Swayamdipta et al. (2020) identify memorized, important, or forgetful training data. However, the notion of data importance in these works is different from how likely the upstream examples will be forgotten during fine-tuning. Furthermore, a systematical study on how such importance is dependent on newly learned tasks is still absent. Prior works represented by Aljundi et al. (2019a); Wang et al. (2024); Aljundi et al. (2019b) study selection strategies of examples for replay-based continual learning algorithms.
Predicting model behaviors. A number of works show LLMs can display a hybrid pattern of unpredictable to highly predictable behaviors (Ganguli et al., 2022; Wei et al., 2022)Ye et al. (2023); Xia et al. (2020); Schram et al. (2023) study prediction of task performance across datasets and training setups. We perform prediction at the example level which is more fine-grained and under-explored.
6 Conclusions
In this paper, we empirically analyzed the associations between learned and forgotten examples in LM fine-tuning. We showed the association displays a low rank pattern across different setups. We showed the example associations alone offer useful information to predict example forgetting when fine-tuning LMs on new tasks. We demonstrated the practical utility of our analysis by showing reduced forgetting as we reweight examples for replay with predicted forgetting. Future works can extend the study to a continual learning setup where new domains or tasks are sequentially learned while predicting forgetting of upstream examples online.
Appendix A Dataset, Model, and LM Training Details
(a) OLMo-7B (LoRA); forgetting over Dolma after full-parameter fine-tuning on FLAN and Tulu.
(b) MPT-7B; forgetting over Dolma after full-parameter fine-tuning on FLAN and Tulu.
(c) OLMo-7B-Instruct (LoRA); forgetting over Tulu after full-parameter fine-tuning on unseen instruction-tuning tasks.
Figure 7: Additional visualized matrices of associations between learned tasks and forgotten examples. We plot forgetting (log-perplexity increase) that occurs on an upstream example (in $x$-axis) after learning a new task (in $y$-axis). Log-perplexity increase can be zero or negative, indicating no forgetting.
Models. We use OLMo-7B [^1] https://huggingface.co/allenai/OLMo-7B of the version pretrained on Dolma v1.6; and OLMo-7B-Instruct [^2] https://huggingface.co/allenai/OLMo-7B-Instruct, which is tuned on Tulu v2 and other human feedback datasets.
(a) Log perplexity increase (forgetting)
(b) Inner products of gradients and weight differences ($z_{ij}^{\textrm{g-w}}$)
(c) Negative inner products of gradients ($z_{ij}^{\textrm{g-g}}$)
Figure 8: A side-by-side comparison between the matrices of forgetting, inner products of gradients and weight differences ($z_{ij}^{\textrm{g-w}}$), and the negative inner products of gradients ($z_{ij}^{\textrm{g-g}}$) we examines in Sec. 3.3.
Learned new tasks and their categorization. We summarize the list and the categorization of newly learned tasks in Tables 3 and 4 in our experiments. We also include the number of training examples and forgetting caused by each task averaged over all upstream examples.
Training and evaluation details. For full-parameter fine-tuning of OLMo-1B and 7B, we train the model for 1,000 steps with an effective batch size of 8 and a linearly decaying learning rate of $2e^{-6}$. For LoRA fine-tuning, we set the rank of adapters as 64 in all our experiments and use a rate of $10^{-4}$. We train the models for 625 steps with an effective batch size of 8. For OLMo-7B-Instruct and MMLU, BBH, TruthfulQA, considering the small size of the training sets, we train the models only for 37 steps with an effective batch size of 8. We use HuggingFace Transformers library for training and VLLM library for efficient inference. The statistics of forgetting are obtained in a single run.
Dataset licenses. MMLU and BBH are released under MIT license. Truthful QA, Dolma, Redpajama, OLMo models, and MPT models are released under Apache 2.0 license. Tulu V2 is released under ODC-By license. Dolly is released under CC BY-SA 3.0 license.
Appendix B Details of Forgetting Prediction and Replay
Data Splits for Predicting Example Forgetting. We mark the tasks used as in-domain test splits for predicting example forgetting (Sec. 4) in Tables 3 and 4. The train-test split for the in-domain tasks is randomly generated.
Training and evaluation details. We use Surprise Library 1.1.3333https://github.com/NicolasHug/Surprise/tree/v1.1.3 for additive linear, SVD, and KNN prediction models. For SVD, we set the dimension of the learnable features as 5. We train the regression models for 100 epochs over the association matrices. The other hyperparameters are left as default.
For in-domain test splits, we randomly sample 30 upstream examples and assume the ground truth forgetting is known for these examples. This is required for predicting forgetting on the rest of upstream examples by additive linear, SVD, and KNN methods. We repeat the experiment 10 times and report the mean and standard deviation in Table 2.
We used OLMo-1B models as the trainable example encoders in the implementation of the prediction method by Jin & Ren (2024) that relies on inner products of trained example representations. At inference, given an upstream example, we compute the averaged dot-product with all examples in the learned task. We note that at inference time the approach does not require ground truth forgetting of a small number of examples. For a fair comparison with other matrix completion methods, we replace the prediction of the approach with ground truth forgetting on these examples.
(a) OLMo-7B (Full)
(b) OLMo-7B (LoRA)
Figure 9: Reconstruction of $Z$ in OLMo-7B experiments with $k$-th singular value and vectors. Higher values of $k$ capture finer-grained details in $Z$.
Replaying upstream examples in fine-tuning. We sparsely replay 1 mini-batch of 8 upstream examples every 32 steps of model update while fine-tuning on new tasks. Given predicted or ground truth forgetting $z_{i,1..J}$ on upstream examples $x_{1..J}$ when learning a new task $T_{i}$, we sample upstream examples to replay from a categorical distribution where $p(x_{j})\propto\exp(z_{i,j}/\tau)$, where $\tau$ is a temperature hyperparameter set as 0.1. The hyperparameter $\tau$ is tuned on a single validation task by using ground truth forgetting $Z$.
Appendix C Additional Results about Example Associations
We visualize progressive reconstruction with $k$-th singular value and singular vectors for OLMo experiments in Figure 9. We see when $k=2$, there is a single row and column with significantly larger forgetting than the others. This pattern exemplifies a complicated association that is not captured by the simple multiplicative model ($k=1$).
Task Category Task Task Category Task
FLAN/Classification aeslc FLAN/QA arc_challenge*
ag_news_subset arc_easy*
imdb_reviews bool_q
sentiment140 coqa*
sst2 cosmos_qa
trec* math_dataset*
yelp_polarity_reviews* natural_questions*
FLAN/Linguistic cola openbookqa*
definite_pronoun_resolution* piqa
fix_punct* trivia_qa*
true_case FLAN/Summarization cnn_dailymail
word_segment gigaword
wsc* multi_news
FLAN/Generation common_gen samsum
copa wiki_lingua_english_en
dart FLAN/Translation para_crawl_enes
e2e_nlg* wmt14_enfr
hellaswag wmt16_translate_csen
opinion_abstracts_idebate* wmt16_translate_deen
opinion_abstracts_rotten_tomatoes wmt16_translate_fien
story_cloze wmt16_translate_roen
web_nlg_en wmt16_translate_ruen*
FLAN/MRC drop wmt16_translate_tren*
multirc Tulu open_orca
quac oasst1
record lima
squad_v1 code_alpaca
squad_v2 gpt4_alpaca
FLAN/NLI anli_r1 cot
anli_r2 science
anli_r3 flan_v2
cb* sharegpt
mnli_matched hard_coded
mnli_mismatched wizardlm
qnli* Dolly brainstorming
rte closed_qa
snli information_extraction
wnli classification
FLAN/Paraphrase glue_mrpc open_qa
glue_qqp* general_qa
paws_wiki creative_writing
stsb summarization
wic*
Table 3: The list of learned tasks in our experiments on OLMo-1B, OLMo-7B and MPT-7B. * notes for tasks used as the in-domain test split in forgetting prediction experiments in Sec. 4.
Task Category Task Task Category Task
MMLU abstract_algebra BBH boolean_expressions*
anatomy causal_judgement
astronomy date_understanding
business_ethics disambiguation_qa
clinical_knowledge dyck_languages*
college_biology* formal_fallacies*
college_chemistry geometric_shapes
college_computer_science hyperbaton*
college_mathematics logical_deduction_five_objects*
college_medicine* logical_deduction_seven_objects
college_physics logical_deduction_three_objects
computer_security movie_recommendation*
conceptual_physics* multistep_arithmetic_two
econometrics navigate
electrical_engineering object_counting*
elementary_mathematics penguins_in_a_table
formal_logic reasoning_about_colored_objects
global_facts* ruin_names
high_school_biology* salient_translation_error_detection
high_school_chemistry snarks
high_school_computer_science sports_understanding
high_school_european_history* temporal_sequences
high_school_geography tracking_shuffled_objects_five_objects
high_school_government_and_politics tracking_shuffled_objects_seven_objects
high_school_macroeconomics tracking_shuffled_objects_three_objects
high_school_mathematics web_of_lies
high_school_microeconomics word_sorting
high_school_physics* TruthfulQA Nutrition
high_school_psychology Stereotypes
high_school_statistics Confusion
high_school_us_history* Psychology
high_school_world_history Language
human_aging* Sociology
human_sexuality* Finance
international_law Indexical Error
jurisprudence Science
logical_fallacies* Misconceptions
machine_learning Economics
management* Education
marketing* Proverbs
medical_genetics Conspiracies
miscellaneous Religion
moral_disputes Statistics
moral_scenarios* Misquotations
nutrition Subjective
philosophy* Law
prehistory History
professional_accounting Fiction
professional_law Mandela Effect
professional_medicine* Politics
professional_psychology Misinformation
public_relations* Logical Falsehood
security_studies Distraction
sociology* Weather
us_foreign_policy* Myths and Fairytales
virology Superstitions
world_religions Advertising
Paranormal
Health
Table 4: The list of learned tasks in our experiments on OLMo-7B-Instruct. * notes for tasks used as the in-domain test split in forgetting prediction experiments in Sec. 4.

### v3

@@
 We demonstrate the practical utility of our analysis by showing statistically significantly reduced forgetting as we upweight predicted examples for replay at fine-tuning.
-Project page: https://inklab.usc.edu/lm-forgetting-prediction/
 1 Introduction
-There has been a growing need for long-term usability of LLMs.
-Although fine-tuning allows efficient and incremental adaptation of models, it risks catastrophic forgetting (McCloskey & Cohen, 1989; Goodfellow et al., 2014) of upstream knowledge learned in the pre-training or instruction-tuning phase, causing unintended prediction changes over known information.
+There has been growing need for promptly updating LLMs to mitigate harmful behaviors, update outdated knowledge, and expand the set of application domains.
+Although fine-tuning allows efficient and incremental updates of models, it risks catastrophic forgetting (McCloskey & Cohen, 1989; Goodfellow et al., 2014) of upstream knowledge learned in the pre-training or instruction-tuning phase, causing unintended prediction changes over known information.
 This is problematic for the stability of online deployed LLM systems, limiting the feasibility of continual fine-tuning in practice (Raffel, 2023; Shi et al., 2024).
@@
 We experiment with OLMo-1B, OLMo-7B, OLMo-7B-Instruct (Groeneveld et al., 2024) and MPT-7B (Team, 2023) models where upstream data is open-source.
-We perform statistics of forgetting (in log perplexity increase) over $N$ upstream examples, after fine-tuning the model on one of $M$ unseen instruction-tuning tasks, and represent the statistics in a $M\times N$ matrix.
+We analyze forgetting (in log perplexity increase) over $N$ upstream examples, after fine-tuning the model on one of $M$ unseen instruction-tuning tasks, and represent the results in a $M\times N$ matrix.
 Afterwards, we visualize the matrices and fit the observations with statistical models to analyze the associations.
@@
 We visualize the matrices reconstructed from each component $s_{k}\bm{\alpha}_{k}\bm{\beta}_{k}^{T}$ in SVD of $Z$.
-Figure 4 provides an example of patterns captured by the $k$-th component.
+Figure 4 provides an example of patterns captured by the $k$-th component in OLMo-7B-Instruct.
 We demonstrate how the visualization extracts patterns of forgetting that is conditional on the learned tasks.
+In Appendix D, we further examine interpretable patterns in the decomposition of $Z$ in OLMo-1B and 7B.
 3.3 How do the Similarity Measures Interpret the Associations?
@@
 Textual similarity.
-We measure textual similarity $z^{\textrm{text}}_{ij}$ between learned tasks and forgotten examples with TF-IDF vectorized over each pair of learned tasks $T_{i}$ and upstream examples $x_{j}$.
+We measure textual cosine similarity $z^{\textrm{text}}_{ij}$ between learned tasks and forgotten examples with TF-IDF vectorized features over each pair of learned tasks $T_{i}$ and upstream examples $x_{j}$.
+We also measure text representation similarity with final layer representations of OLMo-1B.
 Inner products between projected gradients and model weight updates.
@@
 We also measure the negative inner products of the loss gradients between the upstream example $x_{j}$ and a learned task $T_{i}$, $z^{\textrm{g-g}}_{ij}=-\langle\nabla_{\theta}f(x_{j}),\nabla_{\theta}f(T_{i})\rangle$, as an approximation of forgetting (Lopez-Paz & Ranzato, 2017; Chaudhry et al., 2019).
-Table 1: Spearman’s correlation coefficients between various measures of similarity and the actual forgetting on upstream examples after fine-tuning on a task.
-Similarity Metrics Spearman’s $\rho$
-Textual $z^{\textrm{text}}$ 0.004
-Gradient-Weight Difference $z^{\textrm{g-w}}$ 0.005
-Gradient-Gradient $z^{\textrm{g-g}}$0.007
+Table 1: Correlations between various measures of similarity and the actual forgetting on upstream examples after fine-tuning on a task.
+Pearson $\rho$ Spearman $\rho$
+Textual (TF-IDF) -0.049 -0.035
+Textual (Representation) 0.021 0.017
+Gradient-Weight Diff. $z^{\textrm{g-w}}$0.003 -0.009
+Gradient-Gradient Diff. $z^{\textrm{g-g}}$ 0.061 0.052
 Forgetting correlates poorly with similarity measures of learned tasks and upstream examples.
-We evaluate Spearman’s correlation between the actual forgetting $z_{ij}$ and the various similarity measures $\{z^{\textrm{text}}_{ij},z^{\textrm{g-w}}_{ij},z^{\textrm{g-g}}_{ij}\}$ on OLMo-1B and summarize the results in Table 1.
-We notice that none of the similarity measures correlates with the actual forgetting, with a correlation $\rho<0.01$.
+We evaluate correlations between the actual forgetting $z_{ij}$ and the various similarity measures $\{z^{\textrm{text}}_{ij},z^{\textrm{g-w}}_{ij},z^{\textrm{g-g}}_{ij}\}$ on OLMo-1B and summarize the results in Table $1$The table is updated from the earlier preprint (https://arxiv.org/abs/2406.14026v2) due to an implementation issue..
+We notice that none of the similarity measures correlates with the actual forgetting, with a correlation $|\rho|<0.1$.
 An interpretation of the low correlation is that the model weights deviate from the initial weights after fine-tuning to a region where first-order approximation of forgetting does not hold.
@@
 For OLMo-7B-Instruct experiments, we use MMLU and BBH as in-domain tasks and use TruthfulQA and Dolly as out-of-domain testing tasks.
-Details about the tasks included in the training, in-domain testing, and out-of-domain testing sets are discussed in Appendix B.
+Details about the tasks included in the training, in-domain testing, and out-of-domain testing sets are discussed in Tables 7 and 8 in Appendix B.
 Figure 5: The training and testing setup of predicting example forgetting with association matrix completion.
@@
 Matrix completion approaches.
-We run matrix completion algorithms including additive linear, SVD, and k-nearest neighbors (KNN) models.
+We run matrix completion algorithms including additive linear, SVD  [^2] We improved the results of SVD in Table 2 after more hyperparameter tuning in this version., and k-nearest neighbors (KNN) models.
 The additive linear and the SVD models are introduced earlier in Sec.
@@
 Additive 2.81 7.40 3.50 13.33 15.57 6.12 2.81 5.83 7.01 10.02 38.90 21.22
-SVD 2.88 7.53 3.49 12.41 14.98 6.24 2.86 5.84 7.16 8.27 39.03 21.26
+SVD 2.80 7.14 3.48 10.41 13.74 5.89 2.82 5.76 6.80 7.03 40.47 20.23
 KNN 2.79 7.33 3.45 12.80 14.30 5.54 2.84 5.83 6.83 7.71 38.77 20.82
@@
 This is achieved with weighted sampling of upstream examples $x_{j}$ proportional to $\exp{(\hat{z}_{ij}/\tau)}$, where $\hat{z}_{ij}$ are the predicted forgetting and $\tau$ is a temperature hyperparameter set as 0.1.
+As we have discussed in Sec.
+4.1, predicting forgetting with matrix completion requires seed forgetting $z^{S}$ to be evaluated.
+We consider an offline and an online variant of the approach.
+The offline variant performs a replay-free run of fine-tuning on the task $T_{i}$, after which the seed forgetting will be evaluated.
+We then perform another run of fine-tuning while replaying examples with the predicted forgetting.
+This creates computational overhead equivalent to one extra run of fine-tuning, but is still efficient when the training set of fine-tuning is considerably smaller than the upstream data.
+The online variant instead replays random examples for first 10% of fine-tuning steps, after which it evaluates seed forgetting and determine examples to be replayed in the rest of 90% steps.
+Compared to the offline variant, this mitigates the extra overhead of fine-tuning by trading off the prediction accuracy of forgetting.
 Baselines of mitigating forgetting.
 We compare with diverse strategies of selecting upstream examples for sparse replay.
-As a main comparator, we examine whether weighted sampling with the predicted forgetting improves over random sampling of upstream examples (Random).
-We compare with Maximally Interfered Retrieval (MIR) (Aljundi et al., 2019a), a selection strategy sharing the similar notion of importance that forgotten examples should be selected for replay.
+We primarily examine whether weighted sampling with predicted forgetting statistically significantly improves over random sampling of upstream examples (Random).
+We also compare with Maximally Interfered Retrieval (MIR) (Aljundi et al., 2019a), a selection strategy sharing the similar notion of importance that forgotten examples should be selected for replay.
 The approach performs bi-level sampling by selecting the most forgotten examples from a small random subset of upstream data (set as $|\mathcal{S}|=30$).
@@
 3.3, in a similar vein to coreset selection approaches that utilize gradient information (Park et al., 2023; Xia et al., 2024).
+As a reference, we also experiment with upweighting upstream examples with ground truth forgetting $z_{ij}$, which, however, can face computational efficiency issues in practice.
 Metrics.
@@
 We see matrix completion approaches consistently outperform the learned similarity model in the prior work.
-Among the three matrix completion approaches, we notice that KNN models in general achieve the lowest prediction error.
-Given this finding, we primarily leverage predictions given by the KNN model to reduce forgetting during fine-tuning.
+Among the three matrix completion approaches, we notice that SVD models in general achieve the lowest prediction error.
+Besides, KNN in general outperforms additive linear and learned similarity models while being highly computationally efficient.
+Given this finding, we primarily leverage predictions given by the KNN and SVD models to reduce forgetting during fine-tuning.
 (a) 1B, FLAN
@@
 Mitigating forgetting with the predicted forgetting.
-We leverage the forgetting predicted by the KNN model and upweight examples during replay following our procedure in Sec.
+We leverage the online or offline predicted forgetting by the KNN and SVD models to reweight examples during replay following our procedure in Sec.
 4.2.
@@
 We also notice that MIR does not improve over random sampling in our setup, likely because of the small size of the retrieval candidates relative to the upstream examples.
-We see upweighting examples with ground truth forgetting (GT) consistently reduces forgetting compared to random examples.
-Replaying with KNN-predicted forgetting also consistently reduces forgetting, achieving statistical significance in 4 out of 5 tested setups.
-Limitations of replaying with forgetting prediction.
-We note that our current practice of replaying with predicted forgetting requires two runs of fine-tuning: a replay-free run of fine-tuning after which the seed forgetting will be evaluated, and another run while replaying examples with the predicted forgetting.
-This creates computational overhead equivalent to one extra run of fine-tuning, which is identical to the MIR baseline.
-Nevertheless, the practice is still efficient when the training set of fine-tuning is much smaller than the upstream data.
-Future works can predict forgetting on-the-fly during fine-tuning, mitigating the overhead of fine-tuning the model twice.
+Upweighting examples with ground truth forgetting (GT) consistently reduces forgetting compared to random examples.
+By utilizing predicted forgetting by offline SVD and KNN we statistically significantly reduce forgetting compared to random examples in 4 out of 5 tested setups.
+The nuanced differences between SVD and KNN in forgetting mitigation align with their in-domain prediction accuracy of forgetting in Table 2, where SVD outperforms KNN by most on OLMo-7B models.
+Utilizing online-predicted forgetting also statistically significantly improves over replaying random examples in 3 of the setups.
+The gaps between online and offline variants are closer on 7B models than 1B models.
+Effects on downstream task performance measured with task-specific metrics.
+We evaluate fine-tuned OLMo-7B or OLMo-7B-Instruct on unseen LLM leaderboard tasks and present the results in Appendix C.
+We notice the performance stays stable or improves on some downstream tasks while degrades on others, indicating forgetting.
+Although we observe slightly improved performance of Offline-KNN to random or no replay on most forgotten tasks (e.g., Sciq on OLMo-7B and IFEval on OLMo-7B-Instruct), we do not see statistical significance.
+We leave effective algorithms to mitigate downstream task forgetting with predicted forgetting as future work.
+Table 3: Computational cost of replay-based approaches as a summation of fine-tuning costs $FT(\cdot)$, inference costs over upstream examples $EV(\cdot)$, and matrix completion costs $MC$.
+Method Cost
+Random $FT(Y)$
+Ground Truth $2FT(Y)+EV(N)$
+Offline SVD, KNN $2FT(Y)+EV(S)+MC$
+Online SVD, KNN $FT(Y)+EV(S)+MC$
+MIR $2FT(Y)+Y\cdot EV(S)$
+PPL,GradProd $FT(Y)$
+Computational Efficiency.
+Table 3 summarizes the computation cost of the approaches as a function $FT(\cdot)$ of fine-tuning steps, a function $EV(\cdot)$ of upstream examples whose perplexity is evaluated, and the cost of matrix completion (MC) that is much smaller than LLM inference or training.
+We note the total number of upstream examples as $N$, the size of seed examples as $S$, and the number of fine-tuning steps as $Y$.
+As $S$ is much smaller than $M$, the majority of computational costs arise from fine-tuning $FT(Y)$ and $EV(N)$.
+Replaying with ground truth forgetting is the most costly, as it introduces an additional run of fine-tuning (after which forgetting will be evaluated) and inference over potentially very large-scale upstream data.
+As offline prediction of forgetting only mitigates the need for inference, the approach saves computations when the cost of fine-tuning $FT(Y)$ is notably smaller than $EV(N)$, i.e., over small fine-tuning datasets and massive upstream data.
+Online prediction of forgetting is always efficient, requiring only one run of fine-tuning and without the need to evaluate over massive upstream data.
 5 Related Works
@@
 Future works can extend the study to a continual learning setup where new domains or tasks are sequentially learned while predicting forgetting of upstream examples online.
+Limitations.
+In our research process, we tried to control confounding factors such as model size, learning rate, or the number of training steps.
+However, we did not systematically study how the associations between learned tasks and forgotten upstream examples depend on these factors.
+For example, although we presented results of 1B and 7B models from the same family (OLMo) and observe differences in their association statistics (such as $R^{2}$ scores in Sec.
+3.1), it is unclear how the associations will appear in smaller or larger models.
+Besides, our empirical findings raise questions about interpretability of the associations, i.e., why certain examples are more prone to forgetting while learning a new task, and whether there exist more human-interpretable patterns behind the associations.
+Finally, we limited our experiments to fine-tuning on a single task at a time.
+Predicting and mitigating forgetting in sequentially learned tasks is left as future work.
 Appendix A Dataset, Model, and LM Training Details
@@
 Learned new tasks and their categorization.
-We summarize the list and the categorization of newly learned tasks in Tables 3 and 4 in our experiments.
+We summarize the list and the categorization of newly learned tasks in Tables 7 and 8 in our experiments.
 We also include the number of training examples and forgetting caused by each task averaged over all upstream examples.
@@
 We mark the tasks used as in-domain test splits for predicting example forgetting (Sec.
-4) in Tables 3 and 4.
+4) in Tables 7 and 8.
 The train-test split for the in-domain tasks is randomly generated.
 Training and evaluation details.
-We use Surprise Library 1.1.3333https://github.com/NicolasHug/Surprise/tree/v1.1.3 for additive linear, SVD, and KNN prediction models.
+We use Surprise Library 1.1.3555https://github.com/NicolasHug/Surprise/tree/v1.1.3 for additive linear, SVD, and KNN prediction models.
 For SVD, we set the dimension of the learnable features as 5.
@@
 The hyperparameter $\tau$ is tuned on a single validation task by using ground truth forgetting $Z$.
-Appendix C Additional Results about Example Associations
+Appendix C Downstream Task Evaluation with Task-Specific Metrics
+Table 4: Downstream task performance of OLMo-7B models before and after fine-tuning on dolly tasks.
+ARC-Easy ARC-Challenge Boolq Hellaswag Openbookqa Piqa Sciq Winogrande
+Metrics Acc-norm Acc-norm Acc Acc-norm Acc-norm Acc-norm Acc-norm Acc
+Before FT 68.77 40.36 72.41 75.65 42.20 79.54 88.60 66.29
+No Replay 67.34 42.28 74.82 76.89 44.65 80.05 84.09 67.89
+Random 67.48 42.43 74.33 77.26 44.88 79.97 84.77 67.33
+KNN-Offline 67.49 42.24 74.33 77.07 44.30 80.09 84.91 67.54
+Table 5: Downstream task performance of OLMo-7B-Instruct models before and after fine-tuning on dolly tasks.
+MMLU-Pro BBH IF-Eval MUSR GPQA
+Metrics 5-shot Acc 3-shot Acc-norm 0-shot Inst 0-shot Acc-norm 0-shot Acc-norm
+Before FT 18.2 20.27 39.93 20.65 14.34
+No Replay 17.51 20.26 17.72 22.01 14.37
+Random 17.45 20.02 18.19 21.90 14.41
+KNN-Offline 17.45 20.19 18.28 21.90 14.42
+We evaluate downstream task performance of LMs before and after fine-tuning on 8 tasks from Dolly with LM-Evaluation-Harness framework (Gao et al., 2024).
+For OLMo-7B models, we evaluate on the same set of downstream tasks in OLMo technical report (Groeneveld et al., 2024).
+For OLMo-7B-Instruct models, we evaluate on Open LLM Leaderboard tasks [^6] https://huggingface.co/docs/leaderboards/open_llm_leaderboard/about.
+For fine-tuned models, we compare no replay, replaying random examples, and replaying with forgetting predicted by offline KNN.
+Tables 4 and 5 summarize the results.
+We notice that fine-tuning OLMo-7B on Dolly improves downstream performance on most of the downstream tasks.
+This aligns well with the purpose of fine-tuning a LM that is not instruction-tuned.
+Nevertheless, we notice performance degradation on two of the tasks, namely ARC-Easy and Sciq, which indicates forgetting.
+Although offline KNN achieves higher accuracy scores on these two tasks compared to no-replay (67.48 to 67.34, 84.91 to 84.09), we do not find the improvement statistically significant.
+For OLMo-7B-Instruct, fine-tuning on Dolly only improves performance on MUSR.
+The models clearly suffer from forgetting on the other tasks such as IFEval.
+Offline KNN achieves higher scores than random or no replay on IFEval (18.28 compared to 18.19 and 17.72), but we could not conclude about the significance of the improvement.
+To summarize, we do not see clear performance improvement in downstream task performance (evaluated with task-specific metrics) by replaying random or chosen examples.
+We conjecture that replay-based approaches are not sufficient to mitigate forgetting on their own, and can be combined with other approaches such as careful learning rate scheduling or parameter regularization.
+We leave more effective algorithms to mitigate downstream task forgetting with predicted forgetting as future works.
+Appendix D Towards Interpreting Fine-Grained Associations
+Table 6: Semantic meaning of $k$-th component in the SVD of the association matrix $Z$.
+We identify top relevant learned tasks and upstream example domains to $k$-th component in the SVD of the association matrix $Z$.
+OLMo-1B OLMo-7B
+Learned Tasks More Forgotten Domain Learned Tasks More Forgotten Domain
+flan/paws_wiki flan/squad_v2
+$k=1$ flan/glue_mrpc None flan/fix_punct None
+flan/story_cloze tulu/open_orca
+flan/opinion_abstracts_idebate flan/mnli_matched
+$k=2$ dolly/general_qa StackOverflow flan/mnli_mismatched None
+flan/story_cloze flan/snli
+flan/story_cloze flan/squad_v2
+$k=3$ flan/fix_punct None flan/quac None
+flan/true_case flan/fix_punct
+math_dataset flan/rte
+$k=4$ dolly/general_qa None flan/opinion_abstracts_idebate None
+flan/opinion_abstracts_idebate flan/story_cloze
 We visualize progressive reconstruction with $k$-th singular value and singular vectors for OLMo experiments in Figure 9.
-We see when $k=2$, there is a single row and column with significantly larger forgetting than the others.
-This pattern exemplifies a complicated association that is not captured by the simple multiplicative model ($k=1$).
+The visualization exemplifies complicated associations that is not captured by the simple multiplicative model ($k=1$).
+For example, on OLMo-7B (LoRA) and when $k=2$, we see a single row and column with significantly larger forgetting than the others.
+Semantic meanings of $k$-th component in the SVD of the association matrix $Z$.
+We perform further analysis into the patterns captured by the $k$-th singular value and singular vectors by identifying the most relevant learned tasks and upstream example domain to the component.
+For each $k$ and its corresponding component $\hat{Z}_{k}=s_{k}\bm{\alpha}_{k}\bm{\beta}_{k}^{T}$, we extract top 3 rows with the highest mean (i.e., top 3 relevant learned tasks $T_{i}$).
+We also extract top 50 columns with highest mean (i.e.
+top 50 relevant upstream examples) and the domain where these upstream examples are drawn from.
+For OLMo models, the domains are one of C4, common-crawl, Gutenberg books, Reddit, Science, StackOverFlow, and Wikipedia.
+We compare the distribution of domains in the top 50 upstream examples, and perform a $z$-test to determine upstream example domain that is significantly more forgotten compared to a prior domain distribution of top 50 most forgotten upstream examples (colunms with highest mean in $Z$).
+The results are summarized in Table 6.
+We highlight some notable patterns in Table 6.
+(1) Some component $Z_{k}$ highlights forgetting patterns of upstream examples from certain domains.
+On OLMo-1B, the second component ($k=2$) highlights patterns where StackOverFlow examples are forgotten.
+(2) Some component $Z_{k}$ highlight forgetting when learning specific types of tasks.
+For example, the second component ($k=2$) on OLMo-7B highlights forgetting patterns after learning NLI tasks (mnli_matched, mnli_mismatched, snli).
+This also exemplifies how learning similar tasks cause a similar set of upstream examples to be more forgotten.
+We believe a more comprehensive interpretation of the patterns of forgetting is an interesting and challenging future work.
 Task Category Task Task Category Task

### v4

@@
 Large Language models (LLMs) suffer from forgetting of upstream data when fine-tuned.
-Despite efforts on mitigating forgetting, few have investigated whether, and how forgotten upstream examples are dependent on and associated with newly learned tasks.
-Insights on such associations enable efficient and targeted mitigation of forgetting.
-In this paper, we empirically analyze forgetting (measured in log-perplexity increase) that occurs in $N$ upstream examples of language modeling or instruction-tuning after fine-tuning LLMs on one of $M$ new tasks, visualized in $M\times N$ matrices.
-We demonstrate that the matrices display simple low-rank patterns, often well-approximated with multiplicative scalar effects of upstream examples and newly learned tasks.
-We also examine fine-grained associations with visualization and statistics.
-Leveraging the low-rank nature of the associations, we predict forgetting of upstream examples when fine-tuning on unseen tasks with matrix completion over the empirical associations.
+Despite efforts on mitigating forgetting, few have investigated whether, and how forgotten upstream examples are dependent on newly learned tasks.
+Insights on such dependencies enable efficient and targeted mitigation of forgetting.
+In this paper, we empirically analyze forgetting that occurs in $N$ upstream examples of language modeling or instruction-tuning after fine-tuning LLMs on one of $M$ new tasks, visualized in $M\times N$ matrices.
+We show that the matrices are often well-approximated with low-rank matrices, indicating the dominance of simple associations between the learned tasks and forgotten upstream examples.
+Leveraging the analysis, we predict forgetting of upstream examples when fine-tuning on unseen tasks with matrix completion over the empirical associations.
 This enables fast identification of most forgotten examples without expensive inference on the entire upstream data.
@@
 We demonstrate the practical utility of our analysis by showing statistically significantly reduced forgetting as we upweight predicted examples for replay at fine-tuning.
+Catastrophic forgetting; Large language models
 1 Introduction
-There has been growing need for promptly updating LLMs to mitigate harmful behaviors, update outdated knowledge, and expand the set of application domains.
-Although fine-tuning allows efficient and incremental updates of models, it risks catastrophic forgetting (McCloskey & Cohen, 1989; Goodfellow et al., 2014) of upstream knowledge learned in the pre-training or instruction-tuning phase, causing unintended prediction changes over known information.
-This is problematic for the stability of online deployed LLM systems, limiting the feasibility of continual fine-tuning in practice (Raffel, 2023; Shi et al., 2024).
-While extensive works have developed algorithms to mitigate forgetting (Wu et al., 2024), the increasing scale of the model and data necessitates more efficient and targeted approaches.
-A key aspect of the challenge is understanding when and what LLMs forget, prompting analysis into the patterns of frequently forgotten examples (Toneva et al., 2019; Maini et al., 2022; Zhang & Wu, 2024), and the impacts of models and hyperparameters on forgetting (Mirzadeh et al., 2022; Kalajdzievski, 2024; Que et al., 2024; Ibrahim et al., 2024).
-However, how the associations between learned tasks and upstream examples inform forgetting remains under-explored.
-Understanding such associations creates potential for targeted mitigation of forgetting over specific upstream examples when a new task is learned.
-While theoretical and empirical study indicate associations between learned and forgotten tasks in shallower neural networks (Lee et al., 2021; Goldfarb et al., 2024; Ramasesh et al., 2021), the associations are under-explored for LLMs, or measured regarding upstream data of language modeling or instruction-tuning.
-In this paper, we empirically study the associations between learned tasks and forgotten upstream examples of language modeling or instruction-tuning.
-We experiment with OLMo-1B, OLMo-7B, OLMo-7B-Instruct (Groeneveld et al., 2024) and MPT-7B (Team, 2023) models where upstream data is open-source.
-We analyze forgetting (in log perplexity increase) over $N$ upstream examples, after fine-tuning the model on one of $M$ unseen instruction-tuning tasks, and represent the results in a $M\times N$ matrix.
-Afterwards, we visualize the matrices and fit the observations with statistical models to analyze the associations.
-Figure 1 illustrates the analysis setup.
-Our findings suggest that the associations between learned tasks and forgotten examples are usually simple and display a low-rank pattern, as simple regression models can decently fit of the associations.
-We extract more fine-grained associations with visualization and statistics, showing cases where specific upstream examples are forgotten when learning a new task.
-We demonstrate that forgetting does not correlate with many predefined similarity metrics of learned tasks and upstream examples, such as token overlap or gradient dot-products, making these metrics poor predictors of forgetting.
-Inspired by the findings, we directly utilize statistics of forgetting to predict example forgetting on unseen tasks by solving a matrix completion problem over the association matrices, analogical to collaborative filtering (Sarwar et al., 2001) in recommender systems, achieving both efficiency and interpretability.
-Our $k$-nearest neighbor (KNN) model outperforms prior approaches that learn semantic relations of two examples with LMs (Jin & Ren, 2024) .
-We verify the benefit of prediction by upweighting examples with higher predicted forgetting during replay as we fine-tune LLMs on new instruction-tuning tasks, achieving statistically significant improvement in alleviating forgetting compared to replaying random examples.
+There has been growing a need for continued fine-tuning of LLMs to mitigate harmful behaviors, update outdated knowledge, and adapt to unseen tasks and domains.
+Although fine-tuning allows efficient and incremental adaptation of models, models may suffer from catastrophic forgetting (McCloskey & Cohen, 1989; Goodfellow et al., 2014) of upstream knowledge learned in the pre-training or instruction-tuning phase, causing unintended prediction changes over known information.
+This is problematic for the performance and reliability of online deployed LLMs, limiting the feasibility of continual fine-tuning in practice (Raffel, 2023; Shi et al., 2024).
+Existing works demonstrate that replaying or mixing in past examples are effective and scalable approaches to mitigate LLM forgetting (Scialom et al., 2022; Roth et al., 2024; Li et al., 2024a; Ibrahim et al., 2024; Ye et al., 2024).
+These approaches, however, often rely on random sampling of past examples; knowing what models forget after fine-tuning allows more efficient and targeted mitigation of forgetting – e.g., by prioritizing the replay of more forgotten examples (Toneva et al., 2019; Aljundi et al., 2019a).
+In this paper, we explore how forgotten upstream examples depend on the learned tasks, so that forgetting caused by unseen tasks can be efficiently inferred from the forgetting that happened while learning other tasks.
+The complexity of such associations between learned tasks and forgotten examples plays an important role in the predictability; Figure 1 (a) illustrates a hypothetical scenario where certain upstream examples suffer more from forgetting regardless of which tasks are learned, making forgetting easily predictable; in contrast, (b) exemplifies upstream example forgetting that is highly dependent on the learned tasks.
+Existing theoretical and empirical study on the associations between learned and forgotten tasks focus on shallower models (Lee et al., 2021; Goldfarb et al., 2024; Ramasesh et al., 2021); the problem is under-explored for LLM forgetting or in an example level.
+Swayamdipta et al.
+(2020); Maini et al.
+(2022) characterize examples that are prone to forgetting, but they do not touch on how example forgetting depends on the learned tasks.
 Figure 1: The problem setup of analyzing the associations between learned tasks and forgotten upstream examples as we fine-tune LLMs on one of unseen new tasks.
-Over total $N$ upstream examples and $M$ unseen tasks, we measure and record forgetting in a $M\times N$ matrix and analyze the statistical properties of the associations.
-To summarize, the contributions of this paper are (1) an empirical analysis on how forgotten examples are associated with learned tasks in representative 1B and 7B language models, and (2) a novel approach of predicting example forgetting by solving a matrix completion problem over the empirical associations, and (3) a practical and efficient algorithm to mitigate forgetting during LLM fine-tuning by upweighting upstream examples for replay according to the predicted forgetting.
+Over total $N$ upstream examples and $M$ unseen tasks, we measure and record forgetting (in red) in a $M\times N$ matrix and attempt to fit the associations with low-rank approximations.
+Better approximations of low-rank approximations indicate simpler associations between learned tasks and forgotten upstream examples.
+Specifically, we start with analyzing the associations between the learned tasks and forgotten upstream examples in LLM fine-tuning.
+We measure forgetting (in continuous log perplexity increase or binary exact match drop) over $N$ upstream examples, after fine-tuning the model on one of $M$ unseen instruction-tuning tasks, while summarizing the results in a $M\times N$ matrix.
+We evaluate the complexity of the associations by measuring the goodness-of-fit of low-rank approximations of the example associations.
+We then examine how the complexity of the example associations varies across model types (OLMo, OLMo2, MPT, Pythia) and sizes (1B to 13B parameters).
+Our findings suggest that the associations between learned tasks and forgotten examples are often well-approximated with low-rank matrices.
+On OLMo-1B and OLMo-7B, rank-3 approximation fits the associations between 85 learned tasks and 140,000 upstream examples with $R^{2}>0.69$.
+We notice that the forgetting of more capable and recent LLM families are more complicated, requiring higher-rank approximations; within the same model family, the complexity of the associations remains stable or increases with the model size.
+The matrix decomposition further interprets the associations by distinguishing forgetting that are independent of or dependent on what the model learns.
+Following the low-rank approximations of the associations, we predict example forgetting on unseen tasks by solving a matrix completion problem over the association matrices, analogical to collaborative filtering (Sarwar et al., 2001) in recommender systems, achieving both efficiency and interpretability.
+Our matrix factorization (MF) or $k$-nearest neighbor (KNN) models outperform prior approaches that learn semantic relations of two examples with LMs (Jin & Ren, 2024).
+As an example, we achieve 57.35 F1 in predicting binary example forgetting where the F1 of random guess is only 6.3.
+Lastly, we demonstrate the practical benefit of predicting forgetting in mitigating forgetting.
+We upweight upstream examples with higher predicted forgetting during replay as we fine-tune LLMs on new instruction-tuning tasks, achieving statistically significant improvement in alleviating forgetting compared to replaying random examples.
+To summarize, the contributions of this paper are (1) an empirical analysis on how forgotten examples are associated with learned tasks in representative 1B to 13B language models, and (2) a novel approach of predicting example forgetting by solving a matrix completion problem over the empirical associations, and (3) a scalable and efficient algorithm to mitigate forgetting during LLM fine-tuning by upweighting upstream examples for replay according to the predicted forgetting.
 2 Problem and Analysis Setup
-In this section, we start by formally defining the metrics of forgetting and set up the problem formulation of analyzing the associations between learned tasks and forgotten examples.
+In this section, we start by formally defining the metrics of forgetting and set up the problem formulation of analyzing the associations between learned tasks and forgotten upstream examples.
 We then introduce models and datasets used for collecting the statistics.
@@
 Upstream examples and learned tasks.
-Large Language models (LLMs) are commonly pre-trained with language modeling objectives over a massive collection for corpora, with some LLMs further post-trained (instruction-tuned) to better follow human instructions.
+LLMs are commonly pre-trained with language modeling objectives over a massive collection of corpora, and optionally post-trained (instruction-tuned) to better follow human instructions.
 We use upstream data to refer to language modeling or instruction tuning training data used at the pre-training or post-training phase of LLMs.
@@
 We then evaluate performance degradation on each upstream example $x_{i}\in x_{1..N}$.
-We measure log perplexity as the performance metric as they are applicable to both language modeling and instruction tuning, and are known to correlate well with other dataset-specific metrics (Hoffmann et al., 2022).
-For language modeling, log perplexity of an example $x_{i}$ is averaged over all tokens in the example; for instruction tuning, we average log perplexity over the ground truth response tokens only.
-We measure forgetting $z_{ij}$ that occurs on an upstream example $x_{j}\in x_{1..N}$ as increase (degradation) in log perplexity after fine-tuning the LM on a new task $T_{i}\in T_{1..M}$.
+We mainly measure log perplexity as the performance metric as they are applicable to both language modeling and instruction tuning, and are known to correlate well with other dataset-specific metrics (Hoffmann et al., 2022).
+For instruction tuning tasks with a restricted output space (e.g., multi-choice questions), we also measure binary exact matches (EM).
+We measure forgetting $z_{ij}$ that occurs on an upstream example $x_{j}\in x_{1..N}$ as increase in log perplexity or drop of exact match after fine-tuning the LM on a new task $T_{i}\in T_{1..M}$.
 We record forgetting $z_{ij}$ in an association matrix $Z\in\mathbb{R}^{M\times N}$.
-2.2 Models and Datasets
-Our analysis requires access to upstream data of LLMs for measuring their log perplexity changes occurred after fine-tuning.
-We experiment with OLMo-1B, OLMo-7B, OLMo-7B-Instruct, and MPT-7B where upstream data of language modeling and instruction-tuning is open-source.
-OLMo-1B and 7B.
-OLMo models are pre-trained on Dolma (Soldaini et al., 2024), a massive collection of text covering diverse domains such as Wikipedia and news articles.
-We fine-tune LMs over 66 tasks from FLAN-V2 (Longpre et al., 2023), 11 tasks from Tulu V2 (Ivison et al., 2023), and 8 tasks from Dolly (Conover et al., 2023), obtaining 85 fine-tuned models.
-The collections cover diverse instruction-tuning tasks such as reading comprehension, math reasoning, and safety alignment.
-We then evaluate log perplexity increase on a 1% subset of Dolma-v1.6-Sample.
-Each upstream example is a maximum 2,048-token document from Dolma, resulting in 141,816 examples.
-For 7B models, We perform either full-parameter fine-tuning or LoRA adapter tuning (Hu et al., 2022) only, noted as OLMo-7B (Full FT) or OLMo-7B (LoRA) respectively.
-OLMo-7B-Instruct.
-OLMo-7B-Instruct models are instruction-tuned on Tulu V2.
-In our experiments, we few-shot fine-tune OLMo-7B-Instruct over new task data from MMLU (Hendrycks et al., 2021), BBH (Suzgun et al., 2022), TruthfulQA (Lin et al., 2022), and Dolly, and evaluate log perplexity increase over a stratified sample of 10,718 examples from Tulu v2 as upstream examples.
-The number of training examples in these datasets are smaller (on the scale of tens to hundreds of examples), making the setup closer to LLM error editing (Yao et al., 2023; Zhang et al., 2024).
-We perform either full-parameter fine-tuning or LoRA fine-tuning.
-MPT-7B.
-MPT models are pre-trained on a diverse collection of corpora.
-We fine-tune the LM over the same 85 instruction-tuning tasks as OLMo models and evaluate forgetting on Redpajama (Computer, 2023), involved as a part of the pretraining corpora of MPT.
-We sample 10,000 maximum 2,048 token documents from Redpajama as upstream examples.
-We use a learning rate of $2e^{-6}$ for full-parameter fine-tuning and $1e^{-4}$ for LoRA fine-tuning. We include the other training details in Appendix A.
-(a) OLMo-1B; forgetting over Dolma after full-parameter fine-tuning on FLAN and Tulu.
-(b) OLMo-7B; forgetting over Dolma after full-parameter fine-tuning on FLAN and Tulu.
-(c) OLMo-7B-Instruct; forgetting over Tulu after full-parameter fine-tuning on unseen instruction-tuning tasks.
-Figure 2: Visualized matrices of associations between learned tasks and forgotten examples.
+(a) OLMo-7B; forgetting (in red) over Dolma after full-parameter fine-tuning on FLAN, Tulu, and Dolly.
+(b) OLMo-7B-Instruct; forgetting (in red) over Tulu after full-parameter fine-tuning on MMLU, BBH, TruthfulQA, and Dolly.
+Figure 2: Examples of visualized matrices of associations $Z$ between $M$ learned tasks and $N$ upstream examples.
 We plot forgetting (log-perplexity increase) that occurs on an upstream example (in $x$-axis) after learning a new task (in $y$-axis).
 Log-perplexity increase can be zero or negative, which implies no forgetting.
+We visualize binary forgetting measured with exact match drops in Fig.
+10 in Appendix.
+2.2 Models and Datasets
+Our analysis requires open access to upstream data of LLMs.
+We mainly experiment with open-data LLMs including OLMo-1B, OLMo-7B, and OLMo-7B-Instruct (Groeneveld et al., 2024).
+We also include OLMo2-7B, OLMo2-13B (OLMo et al., 2024), MPT-7B (Computer, 2023), and Pythia-1B to 12B (Biderman et al., 2023) for studying the complexity of the example associations across model types and sizes.
+[…]
+Unseen Task $T_{1..M}$ where models are fine-tuned.
+We fine-tune all non-instruction-tuned models over 66 task from FLAN, 11 tasks from Tulu and 8 tasks from Dolly (Conover et al., 2023).
+For OLMo-Instruct, we few-shot fine-tune OLMo-7B-Instruct over new task data from MMLU (Hendrycks et al., 2021), BBH (Suzgun et al., 2022), TruthfulQA (Lin et al., 2022), and Dolly.
+For all models and tasks, we use a learning rate of $2e^{-6}$ for full-parameter fine-tuning and $1e^{-4}$ for LoRA fine-tuning.
+We include the remaining training details in Appendix A.
 3 Associations between Learned Tasks and Forgotten Examples
-In this section, we address the research questions regarding the associations between learned tasks and forgotten upstream examples represented in the $M\times N$ association matrices $Z$.
-We start by examining how complicated are the associations represented by the matrices $Z$ (Sec.
-3.1).
-We then extract and analyze more fine-grained associations in $Z$ (Sec.
-3.2), and examine how the associations correlate with similarity measures of learned tasks and upstream examples (Sec.3.3).
-3.1 How Complicated are the Associations?
-In a hypothetical extreme case, the associations between the learned tasks and forgotten examples can be very simple, e.g., where upstream examples are forgotten regardless of learned tasks; or it can be highly complicated, e.g., where different new tasks cause very specific subsets of upstream examples to be forgotten.
-Figure 1 (right) illustrates how such simple or complicated associations may display in visualized matrices.
-To understand the actual associations, we start by visualizing the association matrix $Z$ collected in the setups described in Sec.
-2.2, and perform quantitative analysis as we examine whether simple regression models can fit $Z$ with low error.
-Analysis on the visualized association matrices.
-We visualize the association matrices $Z$ in Figure 2 for OLMo-1B, OLMo-7B, OLMo-7B-Instruct as we perform full-parameter fine-tuning.
-The matrices of MPT and LoRA fine-tuning of OLMo-7B and OLMo-7B-Instruct are presented in Figure 7 in Appendix.
-We see $Z$ generally displays a neat and simple pattern.
-We notice that certain upstream are more prone to forgetting, while some are never forgotten (displayed as all-zero columns).
-Nevertheless, in some cases, upstream examples that are never forgotten elsewhere are forgotten by learning specific new tasks (e.g.
-wizardlm examples are almost only forgotten after learning tasks from TruthfulQA in Figure 2(c)).
-It implies the association is a mixture of simple and more complicated patterns.
-Quantitative evaluation of simplicity of the associations.
-We quantitatively measure how well the association matrices $Z$ can be approximated with simple regression models with a small number of learnable parameters.
-We consider (1) additive linear models, where $z_{ij}=b+\alpha_{i}+\beta_{j}+\epsilon$, where $\alpha_{i}$ and $\beta_{j}$ are learnable parameters associated with each new task or upstream example.
-(2) multiplicative models (SVD with rank $r$=1), where $z_{ij}=s\alpha_{i}\beta_{j}+\epsilon$.
-Both models involve $M+N$ learnable parameters plus a bias term.
-We then measure $R^{2}$ as the metrics of determining how well the regression models fit the association matrices $Z$.
-Let $f_{ij}$ be the fitted value, $R^{2}$ is defined as $1-\sum_{i,j}(z_{ij}-f_{ij})^{2}/\sum_{i,j}(z_{ij}-\bar{Z})^{2}$.
-$R^{2}$ ranges between 0 and 1, indicating the portion of total variance in $Z$ that is explained by the regression models.
-We report $R^{2}$ of additive and multiplicative models under different setups in Figure 3 (a).
-Multiplicative models fit the associations better.
-In 5 out of 6 setups presented in Figure 3 (a), the multiplicative model achieves a better fit than additive models at the same number of trainable parameters.
-An interpretation of the multiplicative models is that each upstream example is associated with their tendency of being forgotten $(\beta_{j})$; the learned tasks trivially determine how fast all upstream examples are forgotten (with $\alpha_{i}$).
-This also creates cases where certain upstream are never forgotten $(\beta_{j}\approx 0)$, or learn certain tasks causing little forgetting on any upstream examples $(\alpha_{i}\approx 0)$.
-(a) Fitting $Z$ with additive or multiplicative models
-(b) Fitting $Z$ with progressively higher rank matrices
-Figure 3: $R^{2}$ of fitting the associations between learned tasks and forgotten examples with (a) simple additive linear or multiplicative models, or (b) progressively more expressive models.
-A large portion of the variance in the association can be explained by the multiplicative model.
-We notice the multiplicative models achieve $R^{2}$ between 0.43 to 0.72 in different setups despite their simplicity, suggesting generally simple associations between learned tasks and forgotten examples.
-This finding is emphasized by the diversity of the learned tasks we considered (spanning from reading comprehension, coding, to safety alignment) and the broad coverage of domains in upstream language modeling and instruction-tuning data.
-Nevertheless, we also notice the $R^{2}$ scores are relatively lower (0.4 to 0.5) on OLMo-7B and OLMo-7B-Instruct, implying more complicated associations than the other setups (OLMo-1B, MPT-7B, and LoRA fine-tuning of OLMo-7B models), which elicits our next research question.
-3.2 What are More Fine-Grained Associations?
-Figure 4: Patterns in associations captured by $k$-th component in SVD of the assoication matrix $Z$, exemplified with OLMo-7B-Instruct (full-parameter tuning, also in Fig.
-2(c)).
+(a) Fitting Log PPL increase
+(b) Fitting EM Drop
+Figure 3: $R^{2}$ and F1 of low-rank approximations of the example associations with progressively higher rank $r$.
+Forgetting is measured with log perplexity increase or exact match (EM) drop respectively.
+The number of newly learned tasks (a) $M=85$ (b) $M=124$; upstream examples $N>10,000$ in both cases.
+“R” notes for random guess of with the positive class probability in (b).
+In this section, we analyze the associations between learned tasks and forgotten upstream examples represented in the $M\times N$ association matrices $Z$.
+We visualize the association matrix $Z$ collected in the setups described in in Figure 2.
+We formally define low-rank approximations and set up quantitative metrics of the complexity of the associations in Sec.
+3.1, and examine the results of approximation across model types and sizes in Sec.
+3.2.
+Lastly, we try to interpret the extracted associations in Sec.
+3.3.
+3.1 Methods and Metrics of Approximating Example Associations
+To examine whether simple patterns are dominant in the example associations represented by $Z$, we attempt to approximate $Z$ with low rank matrices.
+When $Z$ represents the increase of log perplexity (also the loss of language modeling), we fit matrix factorization models $Z^{r}=\sum_{k=1}^{r}\bm{\alpha}_{k}\bm{\beta}_{k}^{T}$ that minimize the Frobenius norm $||Z-Z^{r}||_{F}$, where $\bm{\alpha}_{k}\in\mathbb{R}^{M}$, $\bm{\beta}_{k}\in\mathbb{R}^{N}$, $r$ is the rank of the matrix decomposition and $k$ is the index of the component.
+For binary forgetting measured with exact match drop, we fit a logistic matrix factorization model $Z^{r}=\sigma(\sum_{k=1}^{r}\bm{\alpha}_{k}\bm{\beta}_{k}^{T})$ that minimizes the cross entropy between $Z^{r}$ and $Z$, where $\sigma$ is the sigmoid function.
+We measure $R^{2}$ or F1-scores of the approximation as the goodness-of-fit metrics.
+Interpretations.
+When $r=1$, the approximation effectively assumes that the forgetting $z_{ij}$ (or its logit) is a scalar product of a parameter $\alpha^{(j)}$ specific to each upstream example and each newly learned task $\beta^{(i)}$.
+The set of more forgotten upstream examples is independent of which task the model learns, as $\beta^{(i)}$ trivially determines how fast forgetting uniformly happens on all upstream examples.
+With a higher rank $r$, the approximation captures task-dependent forgetting where certain upstream examples are disproportionally more forgotten when learning certain tasks.
+This inner product formulation is also connected to the first-order approximation of loss increase as an inner product of weight updates and gradients (Lopez-Paz & Ranzato, 2017; Lee et al., 2019), where $r$ is the number of LLM parameters.
+3.2 Examining Complexity of Example Associations
+(a) $R^{2},r=3$
+(b) $R^{2},r=5$
+(c) Avg. Forgetting
+Figure 4: $R^{2}$ of approximating the example associations $Z$ in $M=19$ Tulu and Dolly tasks and $N=10,000$ randomly sampled upstream examples across different model types and sizes.
+We present reconstruction with rank $r=3$ (a) or $r=5$ (b), and the average forgetting of models in (c).
+General findings.
+We present $R^{2}$ or F1 of fitting the association with $Z^{r}$ with progressively larger rank $r$ in Figure 3.
+We notice that across all training setups of OLMo models, $R^{2}$ quickly increases to 0.69 with $r=3$.
+Notably, even the rank-1 approximation $Z^{1}$ can achieve $R^{2}$ scores higher than 0.5.
+We see similar findings are on binary forgetting measured with OLMo-Instruct models, where F1 can surpass 0.5 with $Z^{1}$, and 0.6 with $Z^{4}$.
+The results suggest that simple patterns are dominant in the associations between learned tasks and forgotten examples.
+Example associations across model types and sizes.
+We compare $R^{2}$ of approximations with a fixed $M$ and $N$ over Pythia, MPT, OLMo, and OLMo2 models.
+Figure 4 summarizes the $R^{2}$ at a given rank $r$.
+We notice that the goodness of approximations differ among model types.
+On Pythia and MPT, $R^{2}$ scores at $Z^{3}$ are higher than 0.88, while on OLMo-7B and OLMo2, the $R^{2}$ scores are around 0.75 and 0.65.
+The size of the models within the same model family also has an impact on $R^{2}$ On Pythia and OLMo2, $R^{2}$ stays stable with slight decrease as the model size increases from 1B to 13B.
+On OLMo, $R^{2}$ is noticeably lower on 7B models compared to 1B.
+Model families that forget more (e.g., Pythia) tend to yield simpler example associations (higher $R^{2}$).
+However, within the same model familiy, OLMo-7B results in a lower $R^{2}$ score forgets than OLMo-1B despite that the average forgetting is higher.
+To summarize, we empirically notice that the associations between learned tasks and forgotten upstream examples are more complicated in more recent and capable LLMs, requiring higher-rank approximations.
+The complexity of the associations stays stable or increases with larger models within the same family.
+In Appendix, we provide more intuitions about how model capability and sizes affect the complexity of the associations with a set of synthetic experiments over MNIST and multi-layer MLPs.
+3.3 Interpreting Example Associations
+Figure 5: Patterns in associations captured by $k$-th component in the matrix factorization of the example associations $Z$, exemplified with OLMo-7B-Instruct (full-parameter tuning).
 The highlighted regions display patterns that TruthfulQA tasks cause more forgetting ($k=2$) and MMLU causes more forgetting on certain tasks in FLAN v2 and Open-Orca ($k=3$).
-We extract more fine-grained associations where fine-tuning on a task causes a specific set of upstream examples to be forgotten.
-We perform Singular Value Decomposition (SVD) over the association matrices $Z$ and progressively reconstruct $Z$ with up to $r$-th singular values and vectors as $Z_{r}=\sum_{k=1}^{r}s_{k}\bm{\alpha}_{k}\bm{\beta}_{k}^{T}$.
-$Z_{r}$ is also the optimal rank-$r$ matrix that minimizes the Frobenius norm $||Z-Z_{r}||_{F}$ and the $R^{2}$ score according to the property of SVD.
-The first component $Z_{1}$ corresponds to the multiplicative model we examined in Sec.
-3.1; the rest of components $Z_{2..M}$ captures fine-grained associations in $Z$.
-SVD reveals fine-grained associations.
-We present $R^{2}$ of fitting the assoication with $Z_{r}$ with progressively larger rank $r$ in Figure 3 (b).
-We notice that for OLMo-7B and OLMo-7B-Instruct (full-parameter fine-tuning), $R^{2}$ quickly increases to 0.69 and 0.78 with $Z_{3}$ (reconstruction with up to the third component).
-We visualize the matrices reconstructed from each component $s_{k}\bm{\alpha}_{k}\bm{\beta}_{k}^{T}$ in SVD of $Z$.
-Figure 4 provides an example of patterns captured by the $k$-th component in OLMo-7B-Instruct.
-We demonstrate how the visualization extracts patterns of forgetting that is conditional on the learned tasks.
-In Appendix D, we further examine interpretable patterns in the decomposition of $Z$ in OLMo-1B and 7B.
-3.3 How do the Similarity Measures Interpret the Associations?
+Patterns of forgetting from matrix factorizations.
+The matrix factorization of $Z$ yields interpretable patterns of forgetting in each of its component $\bm{\alpha}_{k}\bm{\beta}_{k}^{T}$.
+Figure 5 provides an example of patterns captured by the $k$-th component in OLMo-7B-Instruct experiments.
+Similarly, we analyze forgetting on OLMo-1B and 7B and summarize interpretations of each component in Appendix E.
+There are interesting yet semantically intriguing associations.
+For example, on OLMo-7B, Stackoverflow documents are less forgotten when learning summarization tasks, while more forgotten when learning certain paraphrasing tasks.
+Correlations between example associations and similarity measures.
 We examine whether the associations between learned tasks and forgotten examples are interpretable from the similarity between the learned tasks and upstream examples.
-We consider (1) heuristic similarity measures, such as textual similarity, and (2) theoretically grounded approximations, such as inner products of gradients.
-Textual similarity.
-We measure textual cosine similarity $z^{\textrm{text}}_{ij}$ between learned tasks and forgotten examples with TF-IDF vectorized features over each pair of learned tasks $T_{i}$ and upstream examples $x_{j}$.
-We also measure text representation similarity with final layer representations of OLMo-1B.
-Inner products between projected gradients and model weight updates.
-The increase of the log perplexity ${z_{ij}}$ (also known as the cross-entropy loss) can be approximated with inner products $z^{\textrm{g-w}}_{ij}=\langle\nabla_{\theta}f(x_{j}),\theta_{T_{i}}-\theta_{0}\rangle$ under first-order Taylor expansion (Lee et al., 2019; Doan et al., 2020), where $\nabla_{\theta}f(x_{j})$ is the gradient of the loss of $x_{j}$ at the initial model before fine-tuning, and $\theta_{T_{i}}-\theta_{0}$ are the updates in the model weights after fine-tuning.
-Following Park et al.
-(2023); Xia et al.
-(2024), we use a random projection matrix $\bm{P}\sim\mathcal{N}_{|\theta|\times d}(0,1)$ to reduce the dimension of the gradients or the weight changes to save the cost of storing pre-computed statistics, which preserves the inner products with high probability (Johnson & Lindenstrauss, 1984).
-Inner products between projected gradients.
-We also measure the negative inner products of the loss gradients between the upstream example $x_{j}$ and a learned task $T_{i}$, $z^{\textrm{g-g}}_{ij}=-\langle\nabla_{\theta}f(x_{j}),\nabla_{\theta}f(T_{i})\rangle$, as an approximation of forgetting (Lopez-Paz & Ranzato, 2017; Chaudhry et al., 2019).
-Table 1: Correlations between various measures of similarity and the actual forgetting on upstream examples after fine-tuning on a task.
+We consider (1) heuristic similarity measures, such as token or representational similarity, and (2) first-order approximations, such as inner products of gradients and inner products between weight updates and gradients (Lee et al., 2019; Doan et al., 2020).
+We detail each similarity measure in Appendix F.
+We then evaluate correlations between the actual forgetting $z_{ij}$ and the various similarity measures on OLMo-1B and summarize the results in Table 1.
+Forgetting correlates poorly with similarity measures of learned tasks and upstream examples.
+From Table 1, we notice that none of the similarity measures correlates strongly with the actual forgetting, with a correlation $|\rho|<0.1$.
+These results imply that although simple statistical patterns are dominant in the example forgetting, such associations are not well-interpreted with common similarity measures of learned tasks and forgotten examples.
+Therefore, we hypothesize that leveraging the statistics of forgetting allows better prediction of forgetting than the contents of the tasks and examples.
+Table 1: Correlations between various measures of similarity and upstream example forgetting.
 Pearson $\rho$ Spearman $\rho$
@@
 Textual (Representation) 0.021 0.017
-Gradient-Weight Diff. $z^{\textrm{g-w}}$0.003 -0.009
-Gradient-Gradient Diff. $z^{\textrm{g-g}}$ 0.061 0.052
-Forgetting correlates poorly with similarity measures of learned tasks and upstream examples.
-We evaluate correlations between the actual forgetting $z_{ij}$ and the various similarity measures $\{z^{\textrm{text}}_{ij},z^{\textrm{g-w}}_{ij},z^{\textrm{g-g}}_{ij}\}$ on OLMo-1B and summarize the results in Table $1$The table is updated from the earlier preprint (https://arxiv.org/abs/2406.14026v2) due to an implementation issue..
-We notice that none of the similarity measures correlates with the actual forgetting, with a correlation $|\rho|<0.1$.
-An interpretation of the low correlation is that the model weights deviate from the initial weights after fine-tuning to a region where first-order approximation of forgetting does not hold.
-We further visualize the matrices of $z^{\textrm{g-w}}_{ij}$ and $z^{\textrm{g-g}}_{ij}$ in Fig.
-8 in Appendix and provide a side-by-side comparison with the matrices of forgetting $Z$.
-The visualization displays distinct patterns among the three matrices.
-These results imply that although the association matrices $Z$ display a simple pattern, they are not well-interpreted with common similarity measures of learned tasks and forgotten examples.
+[…]
+[…]
 4 Predicting Example Forgetting with Association Matrix Completion
@@
 3 to predict example forgetting as the model learns a new task, a problem also studied in prior works (Jin & Ren, 2024).
-Finding out the most forgotten examples allows better spot of the behavior changes of the models, enabling efficient and targeted approaches to mitigate forgetting, e.g., by replaying these examples (Aljundi et al., 2019a; Wang et al., 2024).
-Although the ground truth forgetting can be directly obtained by running inference with the fine-tuned model over the upstream data, this requires extensive computational resources.
-We restrict the prediction methods to be computationally efficient.
-Our analysis in Sec.
-3 suggests that (1) the associations between learned tasks and forgotten examples display simple statistical patterns, while (2) the associations correlate poorly with many similarity metrics.
-Therefore, we hypothesize that leveraging the statistics of forgetting allows more effective prediction of forgetting than leveraging the contents of the tasks and examples.
-Following this intuition, we formulate prediction of example forgetting as a matrix completion problem over the empirical associations $Z$, analogical to collaborative filtering in recommender systems (Sarwar et al., 2001), where scalable approaches are studied extensively.
-We start by setting up the problem formulation of predicting example forgetting, and evaluate the performance of prediction of different approaches.
-We present reduced forgetting by utilizing the prediction outcomes during fine-tuning.
+Predicting most forgotten examples allows better spot of the behavior changes of the models, enabling efficient and targeted approaches to mitigate forgetting, e.g., by replaying these examples.
+Although the ground truth forgetting can be directly obtained by running inference with the fine-tuned model over the upstream data, this requires extensive computational resources; in constrast, once a prediction model is trained, forgetting caused by unseen tasks can be predicted efficiently.
+Following analysis in Sec.
+3, we formulate prediction of example forgetting as a matrix completion problem over the empirical associations $Z$.
+We start by setting up the problem formulation of predicting example forgetting, and evaluate the performance of different matrix completion algorithms.
+We then demonstrate the practical benefit of predicting example forgetting by utilizing the prediction outcomes to mitigate forgetting during fine-tuning.
 4.1 Training and Evaluation of Forgetting Prediction
@@
 Details about the tasks included in the training, in-domain testing, and out-of-domain testing sets are discussed in Tables 7 and 8 in Appendix B.
-Figure 5: The training and testing setup of predicting example forgetting with association matrix completion.
+Figure 6: The training and testing setup of predicting example forgetting with association matrix completion, and their integration into example replay methods to mitigate forgetting.
+Table 2: RMSE ($\downarrow$) or F1($\uparrow$) of predicting example forgetting over a held-out set of upstream examples after fine-tuning LMs on unseen new tasks.
+We report average performance over different seed sets ($\mathcal{S}$) of upstream examples with known ground truth forgetting beforehand.
+In-Domain Out-of-Domain
+Model OLMo-1B OLMo-7B
+OLMo-7B
+LoRA
+MPT
+OLMo-7B
+Instruct
+OLMo-1B OLMo-7B
+OLMo-7B
+LoRA
+MPT
+OLMo-7B
+Instruct
+Metrics RMSE RMSE RMSE RMSE RMSE F1 RMSE RMSE RMSE RMSE RMSE F1
+Additive 2.81 7.40 3.50 13.33 15.57 54.31 2.81 5.83 7.01 10.02 38.90 43.57
+KNN 2.79 7.33 3.45 12.80 14.30 55.68 2.84 5.83 6.83 7.71 38.77 44.11
+MF 2.80 7.14 3.48 10.41 13.74 57.35 2.82 5.76 6.80 7.03 40.47 42.91
+Similarity 3.84 9.29 5.45 14.00 16.23 55.41 3.93 7.64 8.52 10.97 42.37 42.45
 To apply matrix completion for predicting forgetting, a few entries $z_{ij}$ should be known when a new fine-tuning task $T_{i}\in\mathcal{T}_{\textrm{test}}$ (row $i$) is introduced.
 We therefore assume access to the ground truth forgetting $z_{ij}$ of a tiny random set $\mathcal{S}$ ($|\mathcal{S}|=30$) of upstream examples for $T_{i}\in\mathcal{T}_{\textrm{test}}$, noted as seed forgetting $z_{i}^{\mathcal{S}}=\{z_{ij}|x_{j}\in\mathcal{S}\}$.
-Obtaining seed forgetting typically takes only a few seconds by running inference with the model fine-tuned on $T_{i}$ over $\mathcal{S}$; we then predict forgetting of the rest $10k-100k$ upstream examples.
+Obtaining seed forgetting typically takes only a few seconds by running inference with the model fine-tuned on $T_{i}$ (or fine-tuned for a few steps on $T_{i}$, which we evaluate separately) over $\mathcal{S}$.
+We then predict forgetting of the rest $10k-100k$ upstream examples.
 Figure 6 illustrates an example of the train-test partition, seed forgetting, and the forgetting to be predicted.
-We use Root Mean Squared Error (RMSE) over the $\mathcal{T}_{\textrm{test}}$ as the metrics of predicting example forgetting.
+We use Root Mean Squared Error (RMSE) or F1 over the $\mathcal{T}_{\textrm{test}}$ as the metrics of predicting log-perplexity increase or exact match drop.
 Matrix completion approaches.
-We run matrix completion algorithms including additive linear, SVD  [^2] We improved the results of SVD in Table 2 after more hyperparameter tuning in this version., and k-nearest neighbors (KNN) models.
-The additive linear and the SVD models are introduced earlier in Sec.
+We run matrix completion algorithms including additive linear models, matrix factorization (MF), and k-nearest neighbors (KNN) models.
+The additive linear model approximate forgetting as additive effects of learned tasks $\alpha^{(i)}$ and upstream examples $\beta^{(j)}$ ($\alpha^{(i)}+\beta^{(j)}$).
+The MF models are introduced earlier in Sec.
 3.
@@
 The final prediction is made with a regression head over the inner products of two representations $\langle h(x_{j}),\frac{1}{N_{i}}\sum_{N_{i}}h(x_{i})\rangle$.
-We leave the implementation details of matrix completion approaches and the learned similarity approach in Appendix B.
-Table 2: RMSE of predicting example forgetting over a held-out set of upstream examples after fine-tuning LMs on unseen new tasks.
-We report averaged performance over different seed sets ($\mathcal{S}$) of upstream examples with known ground truth forgetting beforehand.
-In-Domain Out-of-Domain
-OLMo-1B OLMo-7B MPT-7B OLMo-7B-Inst. OLMo-1B OLMo-7B MPT OLMo-7B-Inst.
-Full FT Full FT LoRA Full FT Full FT LoRA Full FT Full FT LoRA Full FT Full FT LoRA
-Additive 2.81 7.40 3.50 13.33 15.57 6.12 2.81 5.83 7.01 10.02 38.90 21.22
-SVD 2.80 7.14 3.48 10.41 13.74 5.89 2.82 5.76 6.80 7.03 40.47 20.23
-KNN 2.79 7.33 3.45 12.80 14.30 5.54 2.84 5.83 6.83 7.71 38.77 20.82
-Similarity 3.84 9.29 5.45 14.00 16.23 6.19 3.93 7.64 8.52 10.97 42.38 23.47
+We leave the implementation details of matrix completion approaches and the learned similarity-based prediction approach in Appendix B.
+(a) 1B, FLAN
+(b) 1B, Tulu&Dolly
+(c) 7B, FLAN
+(d) 7B, Tulu&Dolly
+(e) 7B-Inst, Dolly
+Figure 7: Log perplexity over upstream data with different strategies of selecting examples to be replayed.
+The solid horizontal lines indicate the log perplexity before fine-tuning (i.e., no forgetting).
+The dash lines show the log perplexity achieved by upweighting upstream examples according to their ground truth forgetting.
+* and ** indicate statistical significance of improvement ($p<0.05$ or $p<0.005$) compared to replaying random examples in paired $t$-tests over all fine-tuning tasks.
 4.2 Mitigating Forgetting with Predicted Forgetting
@@
 Compared to the offline variant, this mitigates the extra overhead of fine-tuning by trading off the prediction accuracy of forgetting.
+We illustrate the two variants in Figure 6.
 Baselines of mitigating forgetting.
 We compare with diverse strategies of selecting upstream examples for sparse replay.
-We primarily examine whether weighted sampling with predicted forgetting statistically significantly improves over random sampling of upstream examples (Random).
-We also compare with Maximally Interfered Retrieval (MIR) (Aljundi et al., 2019a), a selection strategy sharing the similar notion of importance that forgotten examples should be selected for replay.
-The approach performs bi-level sampling by selecting the most forgotten examples from a small random subset of upstream data (set as $|\mathcal{S}|=30$).
+We primarily examine whether weighted sampling with predicted forgetting statistically significantly improves over random sampling of upstream examples (Rand).
+We also compare with a variant of Maximally Interfered Retrieval (MIR-T) (Aljundi et al., 2019a), a selection strategy sharing the similar notion of importance that forgotten examples should be selected for replay.
+The approach performs bi-level sampling by selecting the most forgotten examples from a small random subset of upstream data.
+We extend the approach to select forgotten examples after a full training run on a task, instead of single steps, which achieves better performance.
 In addition, we apply strategies that consider different definitions of upstream example importance.
 We examine an approach based on perplexity thresholds (PPL) (Marion et al., 2023), which samples upstream data of which the perplexity is around the median of the distribution.
-For OLMo-1B, we also sample replayed examples proportional to the gradient inner products (Grad-Prod) we evaluated in Sec.
-3.3, in a similar vein to coreset selection approaches that utilize gradient information (Park et al., 2023; Xia et al., 2024).
-As a reference, we also experiment with upweighting upstream examples with ground truth forgetting $z_{ij}$, which, however, can face computational efficiency issues in practice.
+For OLMo-1B, we also sample replayed examples proportional to the gradient inner products (Grad-Prod) (whose correlation with forgetting is evaluated in Table 1), a representative coreset selection approach that utilizes gradient information (Park et al., 2023; Xia et al., 2024).
+As a reference, we also experiment with upweighting upstream examples with ground truth forgetting $z_{ij}$, which, however, is highly inefficient and repetitive to obtain in practice.
 Metrics.
@@
 We see matrix completion approaches consistently outperform the learned similarity model in the prior work.
-Among the three matrix completion approaches, we notice that SVD models in general achieve the lowest prediction error.
+Among the three matrix completion approaches, we notice that MF models in general achieve the lowest prediction error.
 Besides, KNN in general outperforms additive linear and learned similarity models while being highly computationally efficient.
-Given this finding, we primarily leverage predictions given by the KNN and SVD models to reduce forgetting during fine-tuning.
-(a) 1B, FLAN
-(b) 1B, Tulu&Dolly
-(c) 7B, FLAN
-(d) 7B, Tulu&Dolly
-(e) 7B-Inst, Dolly
-Figure 6: Log perplexity over upstream data as replay upstream examples selected by different approaches after fine-tuning OLMo or OLMo-Instruct models over in-domain or out-of-domain test tasks.
-The solid horizontal lines indicate the log perplexity before fine-tuning (i.e., no forgetting).
-The dash lines show the log perplexity achieved by upweighting upstream examples according to their ground truth forgetting.
-* and ** indicate statistical significance of improvement ($p<0.05$ or $p<0.005$) compared to replaying random examples in paired $t$-tests over all fine-tuning tasks.
 Mitigating forgetting with the predicted forgetting.
-We leverage the online or offline predicted forgetting by the KNN and SVD models to reweight examples during replay following our procedure in Sec.
+We leverage the online or offline predicted forgetting by the matrix completion algorithms to reweigh examples during replay following our procedure in Sec.
 4.2.
 Figure 7 summarizes log perplexity after fine-tuning over the held-out (never replayed) upstream data as we apply different upstream example selection approaches.
-We visualize the relative comparisons of these approaches to no forgetting of upstream examples (log perplexity before fine-tuning).
 We notice that example selection based on gradient inner products (Grad) or perplexity threshold (PPL), mainly applied for identifying important training data for a task in prior works, does not show improvement in mitigating forgetting compared to replaying random examples.
 This implies that the notions of example importance in these works are different from how easily the examples are forgotten.
-We also notice that MIR does not improve over random sampling in our setup, likely because of the small size of the retrieval candidates relative to the upstream examples.
+We also notice that MIR-T does not improve over random sampling in our setup, likely because of the small size of the retrieval candidates relative to the upstream examples.
 Upweighting examples with ground truth forgetting (GT) consistently reduces forgetting compared to random examples.
-By utilizing predicted forgetting by offline SVD and KNN we statistically significantly reduce forgetting compared to random examples in 4 out of 5 tested setups.
-The nuanced differences between SVD and KNN in forgetting mitigation align with their in-domain prediction accuracy of forgetting in Table 2, where SVD outperforms KNN by most on OLMo-7B models.
+By utilizing predicted forgetting by offline additive linear, KNN, and MF models, we statistically significantly reduce forgetting compared to random examples in 4 out of 5 tested setups.
+MF models achieves statistical significance in most the scenarios, which aligns with its top average prediction performance of example forgetting in Table 2.
 Utilizing online-predicted forgetting also statistically significantly improves over replaying random examples in 3 of the setups.
 The gaps between online and offline variants are closer on 7B models than 1B models.
+Computational efficiency discussions.
+Table 3 summarizes the computation cost of the approaches as a function $FT(\cdot)$ of fine-tuning steps, a function $EV(\cdot)$ of upstream examples whose perplexity is evaluated, and the cost of matrix completion (MC) that is much smaller than LLM inference or training.
+We note the total number of upstream examples as $N$, the size of seed examples as $S$, and the number of fine-tuning steps as $Y$.
+As $S$ is much smaller than $M$, the majority of computational costs arise from the fine-tuning $FT(Y)$ and the upstream data evaluation $EV(N)$.
+Replaying with ground truth forgetting is the most costly, as it introduces an additional run of fine-tuning (after which forgetting will be evaluated) and inference over potentially very large-scale upstream data.
+The offline prediction and replay approach saves computations in the scenarios of small fine-tuning datasets and massive upstream data, which is often true in practice.
+Online prediction of forgetting does not incur extra cost of fine-tuning or upstream data inference, thereby always being efficient.
 Effects on downstream task performance measured with task-specific metrics.
-We evaluate fine-tuned OLMo-7B or OLMo-7B-Instruct on unseen LLM leaderboard tasks and present the results in Appendix C.
-We notice the performance stays stable or improves on some downstream tasks while degrades on others, indicating forgetting.
-Although we observe slightly improved performance of Offline-KNN to random or no replay on most forgotten tasks (e.g., Sciq on OLMo-7B and IFEval on OLMo-7B-Instruct), we do not see statistical significance.
-We leave effective algorithms to mitigate downstream task forgetting with predicted forgetting as future work.
+We evaluate fine-tuned OLMo-7B or OLMo-7B-Instruct on unseen LLM leaderboard tasks and present the results in Appendix D.
+Although we observe slightly improved performance of replaying predicted examples to random or no replay on most forgotten tasks (e.g., Sciq on OLMo-7B and IFEval on OLMo-7B-Instruct), we do not see statistical significance.
+We leave more effective algorithms to mitigate downstream task forgetting with predicted forgetting as future work, e.g., by integrating predicted forgetting into alternative algorithms that leverage past examples (Aljundi et al., 2017; Buzzega et al., 2020)
 Table 3: Computational cost of replay-based approaches as a summation of fine-tuning costs $FT(\cdot)$, inference costs over upstream examples $EV(\cdot)$, and matrix completion costs $MC$.
+Costs that are minor are displayed in smaller fonts.
 Method Cost
@@
 Ground Truth $2FT(Y)+EV(N)$
-Offline SVD, KNN $2FT(Y)+EV(S)+MC$
-Online SVD, KNN $FT(Y)+EV(S)+MC$
-MIR $2FT(Y)+Y\cdot EV(S)$
+Offline MF $2FT(Y)+\scriptstyle{EV(S)}+\scriptstyle{MC}$
+Online MF $FT(Y)+\scriptstyle{EV(S)}+\scriptstyle{MC}$
+MIR-T $2FT(Y)+Y\cdot\scriptstyle{EV(S)}$
 PPL,GradProd $FT(Y)$
-Computational Efficiency.
-Table 3 summarizes the computation cost of the approaches as a function $FT(\cdot)$ of fine-tuning steps, a function $EV(\cdot)$ of upstream examples whose perplexity is evaluated, and the cost of matrix completion (MC) that is much smaller than LLM inference or training.
-We note the total number of upstream examples as $N$, the size of seed examples as $S$, and the number of fine-tuning steps as $Y$.
-As $S$ is much smaller than $M$, the majority of computational costs arise from fine-tuning $FT(Y)$ and $EV(N)$.
-Replaying with ground truth forgetting is the most costly, as it introduces an additional run of fine-tuning (after which forgetting will be evaluated) and inference over potentially very large-scale upstream data.
-As offline prediction of forgetting only mitigates the need for inference, the approach saves computations when the cost of fine-tuning $FT(Y)$ is notably smaller than $EV(N)$, i.e., over small fine-tuning datasets and massive upstream data.
-Online prediction of forgetting is always efficient, requiring only one run of fine-tuning and without the need to evaluate over massive upstream data.
 5 Related Works
@@
 In this paper, we primarily studied how the associations between learned and forgotten examples inform forgetting.
-Prior works have studied various factors that affect forgetting of the models, such as (1) type and size of the LM (Mehta et al., 2021; Scialom et al., 2022; Kalajdzievski, 2024; Mirzadeh et al., 2022) (2) trainable parts of the model (e.g., LoRA, soft prompts, or full-model tuning) (Biderman et al., 2024a; Razdaibiedina et al., 2023) (3) hyperparameters such as learning rate (Ibrahim et al., 2024; Winata et al., 2023), dropout (Goodfellow et al., 2014), number of training steps (Biderman et al., 2024b; Kleiman et al., 2023) (4) optimizer (Lesort et al., 2023) and training algorithms (e.g., various continual learning algorithms) (Shi et al., 2024; Wu et al., 2024), (5) the upstream examples or the knowledge themselves (Toneva et al., 2019; Zhang & Wu, 2024).
-Future works can study how various factors affect the associations between learned tasks and forgotten examples.
+Prior works have studied various factors that affect forgetting of the models, such as (1) type and size of the LM (Mehta et al., 2021; Scialom et al., 2022; Kalajdzievski, 2024; Mirzadeh et al., 2022) (2) trainable parts of the model (e.g., LoRA, soft prompts, or full-model tuning) (Biderman et al., 2024a; Razdaibiedina et al., 2023) (3) hyperparameters such as learning rate (Ibrahim et al., 2024; Winata et al., 2023), dropout (Goodfellow et al., 2014), number of training steps (Biderman et al., 2024b; Kleiman et al., 2023) (4) optimizer (Lesort et al., 2023) and training algorithms (e.g., various continual learning algorithms) (Smith et al., 2022; Wang et al., 2022; Shi et al., 2024; Wu et al., 2024), (5) the upstream examples or the knowledge themselves (Toneva et al., 2019; Zhang & Wu, 2024).
+Future works can study how these factors affect the associations between learned tasks and forgotten examples.
 We consider empirical and theoretical study on the effect of task similarity on forgetting to be most relevant to ours.
@@
 (2024); Li et al.
-(2024); Liu et al.
+(2024b); Liu et al.
 (2024) study the problem of selecting a subset of training data that maximizes performance on a given domain or task at a fixed budget for LLMs.
@@
 In this paper, we empirically analyzed the associations between learned and forgotten examples in LM fine-tuning.
-We showed the association displays a low rank pattern across different setups.
+We showed that simple low rank patterns are dominant in the example associations and compared the complexity of the associations across model types and sizes.
 We showed the example associations alone offer useful information to predict example forgetting when fine-tuning LMs on new tasks.
-We demonstrated the practical utility of our analysis by showing reduced forgetting as we reweight examples for replay with predicted forgetting.
-Future works can extend the study to a continual learning setup where new domains or tasks are sequentially learned while predicting forgetting of upstream examples online.
-Limitations.
-In our research process, we tried to control confounding factors such as model size, learning rate, or the number of training steps.
-However, we did not systematically study how the associations between learned tasks and forgotten upstream examples depend on these factors.
-For example, although we presented results of 1B and 7B models from the same family (OLMo) and observe differences in their association statistics (such as $R^{2}$ scores in Sec.
-3.1), it is unclear how the associations will appear in smaller or larger models.
-Besides, our empirical findings raise questions about interpretability of the associations, i.e., why certain examples are more prone to forgetting while learning a new task, and whether there exist more human-interpretable patterns behind the associations.
-Finally, we limited our experiments to fine-tuning on a single task at a time.
-Predicting and mitigating forgetting in sequentially learned tasks is left as future work.
+We demonstrated the practical utility of our analysis by showing reduced forgetting as we reweigh examples for replay with predicted forgetting.
+Future works can extend the study to a continual learning setup where new domains or tasks are sequentially.
+Impact Statement
+By analyzing and mitigating forgetting, our research tries to demystify and reduce forgetting in LLM fine-tuning.
+The better understanding alongside the reduced forgetting in turn encourage model developers to promptly update their LLMs to address limitations of the models and improve their models in continuing efforts.
+We expect the broader application of the continual learning practice will reduce training costs compared to re-training models, and ultimately result in more powerful models under a controlled training cost.
 Appendix A Dataset, Model, and LM Training Details
-(a) OLMo-7B (LoRA); forgetting over Dolma after full-parameter fine-tuning on FLAN and Tulu.
-(b) MPT-7B; forgetting over Dolma after full-parameter fine-tuning on FLAN and Tulu.
-(c) OLMo-7B-Instruct (LoRA); forgetting over Tulu after full-parameter fine-tuning on unseen instruction-tuning tasks.
-Figure 7: Additional visualized matrices of associations between learned tasks and forgotten examples.
-We plot forgetting (log-perplexity increase) that occurs on an upstream example (in $x$-axis) after learning a new task (in $y$-axis).
-Log-perplexity increase can be zero or negative, indicating no forgetting.
-Models.
-We use OLMo-7B [^3] https://huggingface.co/allenai/OLMo-7B of the version pretrained on Dolma v1.6; and OLMo-7B-Instruct [^4] https://huggingface.co/allenai/OLMo-7B-Instruct, which is tuned on Tulu v2 and other human feedback datasets.
+Subsample of upstream dataset.
+For OLMo experiments, we sample 141,876 text chunks with length 2,048 from Dolma v1.6-sample as upstream examples.
+For OLMo-7B-Instruct, we randomly sample an approximately balanced number of examples from each task in Tulu, and filter out examples with input length that exceeds 2,048 (the limit of OLMo models) after tokenization.
+This results in 10,718 examples.
+For OLMo2, we sample 70,000 text chunks with length 2,048 from OLMo2-Mix.
+For MPT and Pythia, we sample 10,000 2,048-token text chunks from RedPajama and the Pile respectively.
+Learned new tasks and their categorization. We summarize the list and the categorization of newly learned tasks in Tables 7 and 8 in our experiments.
+Training and evaluation details.
+For full-parameter fine-tuning of non-instruction-tuned LLMs of all types, we train the model for 1,000 steps with an effective batch size of 8 and a linearly decaying learning rate of $2e^{-6}$.
+For LoRA fine-tuning, we set the rank of adapters as 64 in all our experiments and use a rate of $10^{-4}$.
+We train the models for 625 steps with an effective batch size of 8.
+For OLMo-7B-Instruct and MMLU, BBH, TruthfulQA, considering the small size of the training sets, we train the models only for 37 steps with an effective batch size of 8.
+We use HuggingFace Transformers library for training and VLLM library for efficient inference.
+The statistics of forgetting are obtained in a single run.
 (a) Log perplexity increase (forgetting)
@@
 3.3.
-Learned new tasks and their categorization.
-We summarize the list and the categorization of newly learned tasks in Tables 7 and 8 in our experiments.
-We also include the number of training examples and forgetting caused by each task averaged over all upstream examples.
-Training and evaluation details.
-For full-parameter fine-tuning of OLMo-1B and 7B, we train the model for 1,000 steps with an effective batch size of 8 and a linearly decaying learning rate of $2e^{-6}$.
-For LoRA fine-tuning, we set the rank of adapters as 64 in all our experiments and use a rate of $10^{-4}$.
-We train the models for 625 steps with an effective batch size of 8.
-For OLMo-7B-Instruct and MMLU, BBH, TruthfulQA, considering the small size of the training sets, we train the models only for 37 steps with an effective batch size of 8.
-We use HuggingFace Transformers library for training and VLLM library for efficient inference.
-The statistics of forgetting are obtained in a single run.
-Dataset licenses.
-MMLU and BBH are released under MIT license.
-Truthful QA, Dolma, Redpajama, OLMo models, and MPT models are released under Apache 2.0 license.
-Tulu V2 is released under ODC-By license.
-Dolly is released under CC BY-SA 3.0 license.
+(a) OLMo-7B-Instruct; binary forgetting over FLAN after full-parameter fine-tuning on unseen instruction-tuning tasks.
+(b) OLMo-7B-Instruct (LoRA); binary forgetting over FLAN after full-parameter fine-tuning on unseen instruction-tuning tasks.
+Figure 10: Visualized matrices of associations between learned tasks and forgotten examples measured with binary exact match drops.
 Appendix B Details of Forgetting Prediction and Replay
@@
 Training and evaluation details.
-We use Surprise Library 1.1.3555https://github.com/NicolasHug/Surprise/tree/v1.1.3 for additive linear, SVD, and KNN prediction models.
-For SVD, we set the dimension of the learnable features as 5.
-We train the regression models for 100 epochs over the association matrices.
-The other hyperparameters are left as default.
+We use Surprise Library 1.1.3111https://github.com/NicolasHug/Surprise/tree/v1.1.3 for additive linear, MF, and KNN prediction models.
+For MF, we set the dimension of the learnable features as 5.
+We train the regression models for 1,000 epochs over the association matrices.
 For in-domain test splits, we randomly sample 30 upstream examples and assume the ground truth forgetting is known for these examples.
-This is required for predicting forgetting on the rest of upstream examples by additive linear, SVD, and KNN methods.
-We repeat the experiment 10 times and report the mean and standard deviation in Table 2.
-We used OLMo-1B models as the trainable example encoders in the implementation of the prediction method by Jin & Ren (2024) that relies on inner products of trained example representations.
+This is required for predicting forgetting on the rest of upstream examples by additive linear, MF, and KNN methods.
+We repeat the experiment 10 times.
+Table 4: Downstream task performance of OLMo-7B models before and after fine-tuning on dolly tasks.
+ARC-Easy ARC-Challenge Boolq Hellaswag Openbookqa Piqa Sciq Winogrande
+Metrics Acc-norm Acc-norm Acc Acc-norm Acc-norm Acc-norm Acc-norm Acc
+Before FT 68.77 40.36 72.41 75.65 42.20 79.54 88.60 66.29
+No Replay 67.34 42.28 74.82 76.89 44.65 80.05 84.09 67.89
+Random 67.48 42.43 74.33 77.26 44.88 79.97 84.77 67.33
+KNN-Offline 67.49 42.24 74.33 77.07 44.30 80.09 84.91 67.54
+We used OLMo-1B models as the trainable example encoders in the implementation of the prediction method by (Jin & Ren, 2024) that relies on inner products of trained example representations.
 At inference, given an upstream example, we compute the averaged dot-product with all examples in the learned task.
@@
 The hyperparameter $\tau$ is tuned on a single validation task by using ground truth forgetting $Z$.
-Appendix C Downstream Task Evaluation with Task-Specific Metrics
-Table 4: Downstream task performance of OLMo-7B models before and after fine-tuning on dolly tasks.
-ARC-Easy ARC-Challenge Boolq Hellaswag Openbookqa Piqa Sciq Winogrande
-Metrics Acc-norm Acc-norm Acc Acc-norm Acc-norm Acc-norm Acc-norm Acc
-Before FT 68.77 40.36 72.41 75.65 42.20 79.54 88.60 66.29
-No Replay 67.34 42.28 74.82 76.89 44.65 80.05 84.09 67.89
-Random 67.48 42.43 74.33 77.26 44.88 79.97 84.77 67.33
-KNN-Offline 67.49 42.24 74.33 77.07 44.30 80.09 84.91 67.54
-Table 5: Downstream task performance of OLMo-7B-Instruct models before and after fine-tuning on dolly tasks.
-MMLU-Pro BBH IF-Eval MUSR GPQA
-Metrics 5-shot Acc 3-shot Acc-norm 0-shot Inst 0-shot Acc-norm 0-shot Acc-norm
-Before FT 18.2 20.27 39.93 20.65 14.34
-No Replay 17.51 20.26 17.72 22.01 14.37
-Random 17.45 20.02 18.19 21.90 14.41
-KNN-Offline 17.45 20.19 18.28 21.90 14.42
+Appendix C Details of the Example Similarity Metrics
+In this section, we detail the example similarity metrics applied in our analysis in Sec. 3.3.
+Textual similarity.
+We measure textual cosine similarity $z^{\textrm{text}}_{ij}$ between learned tasks and forgotten examples with TF-IDF vectorized features over each pair of learned tasks $T_{i}$ and upstream examples $x_{j}$.
+We also measure text representation similarity with final layer representations of OLMo-1B.
+Inner products between projected gradients and model weight updates.
+The increase of the log perplexity ${z_{ij}}$ can be approximated with inner products $z^{\textrm{g-w}}_{ij}=\langle\nabla_{\theta}f(x_{j}),\theta_{T_{i}}-\theta_{0}\rangle$ under first-order Taylor expansion (Lee et al., 2019; Doan et al., 2020), where $\nabla_{\theta}f(x_{j})$ is the gradient of the loss of $x_{j}$ at the initial model before fine-tuning, and $\theta_{T_{i}}-\theta_{0}$ are the updates in the model weights after fine-tuning.
+Following (Park et al., 2023; Xia et al., 2024), we use a random projection matrix $\bm{P}\sim\mathcal{N}_{|\theta|\times d}(0,1)$ to reduce the dimension of the gradients or the weight changes to save the cost of storing pre-computed statistics, which preserves the inner products with high probability (Johnson & Lindenstrauss, 1984).
+Inner products between projected gradients.
+We also measure the negative inner products of the loss gradients between the upstream example $x_{j}$ and a learned task $T_{i}$, $z^{\textrm{g-g}}_{ij}=-\langle\nabla_{\theta}f(x_{j}),\nabla_{\theta}f(T_{i})\rangle$, as an approximation of forgetting (Lopez-Paz & Ranzato, 2017; Chaudhry et al., 2019).
+Appendix D Downstream Task Evaluation with Task-Specific Metrics
 We evaluate downstream task performance of LMs before and after fine-tuning on 8 tasks from Dolly with LM-Evaluation-Harness framework (Gao et al., 2024).
@@
 Offline KNN achieves higher scores than random or no replay on IFEval (18.28 compared to 18.19 and 17.72), but we could not conclude about the significance of the improvement.
-To summarize, we do not see clear performance improvement in downstream task performance (evaluated with task-specific metrics) by replaying random or chosen examples.
+To summarize, we do not see clearly reduced forgetting in downstream task performance (evaluated with task-specific metrics) by replaying random or chosen examples.
 We conjecture that replay-based approaches are not sufficient to mitigate forgetting on their own, and can be combined with other approaches such as careful learning rate scheduling or parameter regularization.
 We leave more effective algorithms to mitigate downstream task forgetting with predicted forgetting as future works.
-Appendix D Towards Interpreting Fine-Grained Associations
-Table 6: Semantic meaning of $k$-th component in the SVD of the association matrix $Z$.
-We identify top relevant learned tasks and upstream example domains to $k$-th component in the SVD of the association matrix $Z$.
-OLMo-1B OLMo-7B
-Learned Tasks More Forgotten Domain Learned Tasks More Forgotten Domain
+Table 5: Downstream task performance of OLMo-7B-Instruct models before and after fine-tuning on dolly tasks.
+MMLU-Pro BBH IF-Eval MUSR GPQA
+Metrics 5-shot Acc 3-shot Acc-norm 0-shot Inst 0-shot Acc-norm 0-shot Acc-norm
+Before FT 18.20 20.27 39.93 20.65 14.34
+No Replay 17.51 20.26 17.72 22.01 14.37
+Random 17.45 20.02 18.19 21.90 14.41
+KNN-Offline 17.45 20.19 18.28 21.90 14.42
+Table 6: Semantic meaning of $k$-th component in the factorization of the example association matrix $Z$.
+We identify top relevant learned tasks and upstream example domains to $k$-th component in the factorization of the example association matrix $Z$.
+OLMo-7B OLMo-1B
+Learned Tasks More(+) / Less(-) Forgotten Domain Learned Tasks More(+) / Less (-) Forgotten Domain
 flan/paws_wiki flan/squad_v2
@@
 flan/opinion_abstracts_idebate flan/mnli_matched
-$k=2$ dolly/general_qa StackOverflow flan/mnli_mismatched None
+$k=2$ dolly/general_qa StackOverflow (-) flan/mnli_mismatched None
 flan/story_cloze flan/snli
@@
 flan/opinion_abstracts_idebate flan/story_cloze
+Appendix E Towards Interpreting Fine-Grained Associations
 We visualize progressive reconstruction with $k$-th singular value and singular vectors for OLMo experiments in Figure 11.
@@
 For example, on OLMo-7B (LoRA) and when $k=2$, we see a single row and column with significantly larger forgetting than the others.
-Semantic meanings of $k$-th component in the SVD of the association matrix $Z$.
+Semantic meanings of $k$-th component in the low-rank approximation of the association matrix $Z$.
 We perform further analysis into the patterns captured by the $k$-th singular value and singular vectors by identifying the most relevant learned tasks and upstream example domain to the component.
-For each $k$ and its corresponding component $\hat{Z}_{k}=s_{k}\bm{\alpha}_{k}\bm{\beta}_{k}^{T}$, we extract top 3 rows with the highest mean (i.e., top 3 relevant learned tasks $T_{i}$).
+For each $k$ and its corresponding component $\bm{\alpha}_{k}\bm{\beta}_{k}^{T}$, we extract top 3 rows with the highest mean (i.e., top 3 relevant learned tasks $T_{i}$).
 We also extract top 50 columns with highest mean (i.e.
@@
 For OLMo models, the domains are one of C4, common-crawl, Gutenberg books, Reddit, Science, StackOverFlow, and Wikipedia.
-We compare the distribution of domains in the top 50 upstream examples, and perform a $z$-test to determine upstream example domain that is significantly more forgotten compared to a prior domain distribution of top 50 most forgotten upstream examples (colunms with highest mean in $Z$).
+We compare the distribution of domains in the top 50 upstream examples, and perform a $z$-test to determine upstream example domain that is significantly more or less forgotten compared to a prior domain distribution of top 50 most forgotten upstream examples (columns with highest mean in $Z$).
 The results are summarized in Table 6.
@@
 (1) Some component $Z_{k}$ highlights forgetting patterns of upstream examples from certain domains.
-On OLMo-1B, the second component ($k=2$) highlights patterns where StackOverFlow examples are forgotten.
+On OLMo-7B, the second component ($k=2$) highlights patterns where StackOverFlow examples are disproportionally less forgotten.
 (2) Some component $Z_{k}$ highlight forgetting when learning specific types of tasks.
-For example, the second component ($k=2$) on OLMo-7B highlights forgetting patterns after learning NLI tasks (mnli_matched, mnli_mismatched, snli).
+For example, the second component ($k=2$) on OLMo-1B highlights forgetting patterns after learning NLI tasks (mnli_matched, mnli_mismatched, snli).
 This also exemplifies how learning similar tasks cause a similar set of upstream examples to be more forgotten.
-We believe a more comprehensive interpretation of the patterns of forgetting is an interesting and challenging future work.
+(a) Overlapping rotations
+(b) Disjoint rotations
+(c) Avg. Forgetting
+Figure 12: $R^{2}$ of low-rank approximations of the example association in the forgetting of MLP models on rotated MNIST experiments.
+We report $R^{2}$ when the rotation of the newly learned task overlaps or is disjoint with the upstream data separately.
+$y$-axes are not in the same scale.
+Appendix F Synthetic Dataset Experiments
+In this section, we present a set of experiments on a synthetic dataset, Rotated-MNIST, broadly used in continual learning research.
+We aim to provide intuition for future research about how the complexity of the example associations can depend on (1) the coverage of knowledge represented in upstream examples, and (2) the size of the models, in highly controlled setups.
+We apply a training setup that resembles pretraining and fine-tuning paradigm in transformer language models.
+Specifically, the models are first pre-trained on 10 rotated variants (0-90) of the MNIST digit classification dataset (as upstream tasks and examples).
+Then, the models are fine-tuned on one of 40 unseen rotations for one epoch.
+Among the 40 unseen rotations, 20 are drawn from the same range as upstream examples (0-90), while the other 20 are drawn from a disjoint range (-90-0).
+This separation controls the amount of shared knowledge between the newly learned and upstream tasks.
+We mostly apply other training setups and hyperparameters in Aljundi et al.
+(2019a).
+Each task (rotation) includes 1,000 training examples.
+We train a MLP classifier to predict among 10 digits given an input image without providing its rotation (or the task identifier).
+We experiment MLP classification models with 1 layer (a linear model) to 5 layers.
+We collect the example associations $Z$, and visualize them in Figures 13 and 14.
+The upstream and the newly learned rotation tasks are ordered by their rotations in the $x$ or $y$ axis.
+(a) Linear
+(b) 2-Layer MLP
+(c) 3-Layer MLP
+(d) 4-Layer MLP
+(e) 5-Layer MLP
+Figure 13: Example associations between learned tasks and forgotten upstream examples on Rotated MNIST with overlapping ranges of rotations.
+We measure increase in the cross-entropy loss as forgetting.
+(a) Linear
+(b) 2-Layer MLP
+(c) 3-Layer MLP
+(d) 4-Layer MLP
+(e) 5-Layer MLP
+Figure 14: Example associations between learned tasks and forgotten upstream examples on Rotated MNIST with disjoint ranges of rotations.
+Effects of knowledge coverage.
+When the rotations of the newly learned tasks overlaps with the range of upstream examples, the forgetting is harder to approximate with low-rank approximations, resulting in a $R^{2}$ only around 0.5 for all MLP models with rank-1 approximation.
+In contrast, the $R^{2}$ scores are much higher when the rotations do not overlap, alongside higher average forgetting; $R^{2}$ with rank-1 approximation is higher than 0.8 in these setups.
+The results imply that the amount of shared knowledge between fine-tuning tasks and upstream examples can have an impact on the complexity of the example associations.
+In other words, models trained with a broad coverage of upstream examples, or that cover knowledge required for diverse fine-tuning tasks, can yield more complicated example associations.
+In the context of LLMs, we have noticed that more powerful LLMs (such as OLMo and OLMo2, compared to Pythia and MPT) with broader coverage of knowledge yield more complicated patterns of forgetting.
+Effects of model sizes.
+We compare the MLPs of different number of layers trained on the same upstream data.
+From Figure 12, the patterns of forgetting are nosier in deeper models.
+The $R^{2}$ scores at rank 3 or 5 decreases with added layers in MLPs, providing a quantitative measure of increased complexity between learned tasks and forgotten examples.
+To summarize, our analysis with synthetic datasets provide intuition about the effect of knowledge coverage and model sizes on the complexity of example associations in forgetting.
+We hope the set of synthetic experiments can inspire more comprehensive study on how the complexity of example associations are affected by various factors in future works.
 Task Category Task Task Category Task